### v2

# Asymptotic and compound e-values: multiple testing and empirical Bayes

Abstract
We explicitly define the notions of (exact, approximate or asymptotic) compound p-values and e-values, which have been implicitly presented and extensively used in the recent multiple testing literature. While it is known that the e-BH procedure with compound e-values controls the FDR, we show the converse: every FDR controlling procedure can be recovered by instantiating the e-BH procedure with certain compound e-values. Since compound e-values are closed under averaging, this allows for combination and derandomization of FDR procedures. We then connect compound e-values to empirical Bayes. In particular, we use the fundamental theorem of compound decision theory to derive the log-optimal simple separable compound e-value for testing a set of point nulls against point alternatives: it is a ratio of mixture likelihoods. We extend universal inference to the compound setting. As one example, we construct approximate compound e-values for multiple t-tests, where the (nuisance) variances may be different across hypotheses. Finally, we provide connections to related notions in the literature stated in terms of p-values.
1 Introduction
The main purpose of this paper is to formalize the notions of asymptotic e-values, compound e-values and asymptotic compound e-values, show their versatility and usefulness for multiple testing, as well as provide certain optimal constructions and examples.
Contributions and Paper Outline.
In Section 2, we formally define approximate p-values and e-values, which are the building blocks of all contents in our paper, and provide several sets of conditions that lead to approximate p-values and e-values. In particular, Theorem 2.3 shows that calibration between exact p-values and exact e-values remains valid for approximate p-values and e-values, preserving their approximation parameters. This result will be useful in several constructions later in the paper. In Section 3, we introduce asymptotic p-values and e-values, connect them with usual limiting results in asymptotic statistics, and give a prototypical example of asymptotic e-values. In Section 4, we move to the setting of multiple testing, by defining compound e-values, studying their properties, and giving several examples. Their approximate and asymptotic versions are then discussed in Section 5, and a prototypical construction of approximate compound e-values for multiple t-tests is provided in Section 5.3. In Section 6, we describe the fundamental role of compound e-values in multiple testing: Theorem 6.2 (in Section 6.1) shows how to construct compound e-values from any FDR-controlling procedure, while Theorem 6.5 (in Section 6.2) demonstrates that every FDR procedure can be recovered by e-BH applied to compound e-values. We explain how this framework enables combination and derandomization of multiple testing procedures (Section 6.3) and that compound e-values provide a natural notion of weights for the weighted p-BH procedure (Section 6.4). In Section 7, we study sequence models and develop connections to compound decision theory, which motivates the nomenclature “compound e-values:” In Section 7.1, we provide a bird’s eye overview of compound decision theory, while in Section 7.2, we construct log-optimal simple separable compound e-values (Theorem 7.2 provides a formal statement). In Section 7.3, we introduce compound universal inference (CUI) and empirical Bayes universal inference (EBUI) for composite null hypotheses, proving in Proposition 7.4 that these constructions dominate universal inference e-values.
2 Approximate p-values and e-values
We first review the definition of e-variables and p-variables in single hypothesis testing. Suppose we observe data $X$ drawn according to some unknown probability distribution $\mathbb{P}^{*}$. (Here, $X$ denotes the entire dataset available to the researcher.) The underlying sample space $\Omega$ is left implicit. We define a hypothesis (null or alternative) as a set of probability measures. We reserve the notation $\mathcal{P}$ for a set of probability measures when referring to a null hypothesis. That is, we define a null hypothesis as $H:\mathbb{P}^{*}\in\mathcal{P}$ and say in words that $\mathcal{P}$ contains the unknown true data generating distribution. Analogously, we reserve the notation $\mathcal{Q}$ for the set of probability measures that specify the alternative hypothesis. A p-variable $P=P(X)$ for hypothesis $H$ is a nonnegative random variable that satisfies $\mathbb{P}(P\leq t)\leq t$ for all $t\in[0,1]$ and all $\mathbb{P}\in\mathcal{P}$. Meanwhile, an e-variable $E=E(X)$ for hypothesis $H$ is a $[0,\infty]$-valued random variable satisfying $\mathbb{E}^{\mathbb{P}}[E]\leq 1$ for all $\mathbb{P}\in\mathcal{P}$. The realizations of e- or p-variables are called e- or p-values. As is common in the literature, we use the terms e-values (or p-values) to refer to both the random variables and their realizations; the intended meaning should be clear from the context.
Let $\mathbb{R}_{+}=[0,\infty)$. We say that a probability $\mathbb{P}$ is atomless if there exists a uniformly distributed random variable on $[0,1]$ under $\mathbb{P}$. In what follows, $x\wedge y=\min\{x,y\}$ and $x\vee y=\max\{x,y\}$.
Definition 2.1 (Approximate p-variables and e-variables).
Fix the null hypothesis $\mathcal{P}$ and two functions $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$, and write $\varepsilon_{\mathbb{P}}=\varepsilon(\mathbb{P})$ and $\delta_{\mathbb{P}}=\delta(\mathbb{P})$ for $\mathbb{P}\in\mathcal{P}$.
  1. (i)
    A nonnegative random variable $P$ is an $(\varepsilon,\delta)$-approximate p-variable for $\mathcal{P}$ if
    $\mathbb{P}(P\leq t)\leq(1+\varepsilon_{\mathbb{P}})t+\delta_{\mathbb{P}}\qquad\mbox{for all $t\in(0,1)$ and all $\mathbb{P}\in\mathcal{P}$.}$
  2. (ii)
    A $[0,\infty]$-valued random variable $E$ is an $(\varepsilon,\delta)$-approximate e-variable for $\mathcal{P}$ if
    $\mathbb{E}^{\mathbb{P}}[E\wedge t]\leq 1+\varepsilon_{\mathbb{P}}+\delta_{\mathbb{P}}t\qquad\mbox{for all $t\in\mathbb{R}_{+}$ and all $\mathbb{P}\in\mathcal{P}$.}$
For two constants $\varepsilon\in\mathbb{R}_{+},$ and $\delta\in[0,1)$, $(\varepsilon,\delta)$-approximation is understood as those in Definition 2.1 with the corresponding constant functions.
Wang and Ramdas (2022, Section 6.2) considered $(\varepsilon,0)$-approximate e-variables (not using this terminology).
We make some straightforward observations. Clearly, $(0,0)$-approximate p-variables (e-variables) are just p-variables (e-variables). For a constant $\varepsilon\in\mathbb{R}_{+}$, $P$ is an $(\varepsilon,0)$-approximate p-variable if and only if $(1+\varepsilon)P$ is a p-variable, and $E$ is an $(\varepsilon,0)$-approximate e-variable if and only if $E/(1+\varepsilon)$ is an e-variable. The $\delta$-parameter is more lenient than the $\varepsilon$-parameter; for instance, $(\varepsilon,\delta)$-approximation implies $(0,\varepsilon+\delta)$-approximation in both classes (see Proposition 2.5 below for a sharper result). If $\delta=1$, then $(\varepsilon,\delta)$-approximation puts no requirements on $P$ or $E$.
The definitions of approximate p-variables and approximate e-variables can be interpreted as allowing for a wiggle room in the definitions of p-variables and e-variables in two ways: (a) the probability and expectation requirements have a multiplicative factor $1+\varepsilon_{\mathbb{P}}$; (b) the requirements are only required to hold on an event of probability $1-\delta_{\mathbb{P}}$. This will be made transparent through the following proposition.
Proposition 2.2.
Let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$ and $P$ and $E$ be nonnegative random variables ($[0,\infty]$-valued for $E$). For the following statements,
  1. (P1)
    $P$ is an $(\varepsilon,\delta)$-approximate p-variable for $\mathcal{P}$;
  2. (P2)
    for every $\mathbb{P}\in\mathcal{P}$, there exists an event $A$ such that
    $\mathbb{P}(P\leq t,A)\leq(1+\varepsilon_{\mathbb{P}})t~{}\mbox{for all $t\in(0,1)$ ~{}~{}and~{}~{} }\mathbb{P}(A)\geq 1-\delta_{\mathbb{P}};$
  3. (E1)
    $E$ is an $(\varepsilon,\delta)$-approximate e-variable for $\mathcal{P}$;
  4. (E2)
    for every $\mathbb{P}\in\mathcal{P}$, there exists an event $A$ such that
    $\mathbb{E}^{\mathbb{P}}[E\mathds{1}_{A}]\leq 1+\varepsilon_{\mathbb{P}}\mbox{ ~{}~{}and~{}~{} }\mathbb{P}(A)\geq 1-\delta_{\mathbb{P}};$
we have that (P2) implies (P1) and (E2) implies (E1). Moreover, if each $\mathbb{P}\in\mathcal{P}$ is atomless, then (P1)–(P2) are equivalent, and (E1)–(E2) are equivalent.
Proof.
When $\mathbb{P}$ is atomless, for any random variable $X$, there exists an event $A$ with $\mathbb{P}(A)=1-\delta_{\mathbb{P}}$ such that $\{X<t\}\subseteq A\subseteq\{X\leq t\}$ for some $t\in\mathbb{R}$; see Lemma A.3 of Wang and Zitikis (2021). This fact will be used in the proof below. If $\delta_{\mathbb{P}}=1$ there is nothing to prove, so we assume $\delta_{\mathbb{P}}<1$ below.
(P2) $\Rightarrow$ (P1): It suffices to notice
$\mathbb{P}(P\leq t)\leq\mathbb{P}(P\leq t,A)+\mathbb{P}(A^{c})\leq(1+\varepsilon_{\mathbb{P}})t+\delta_{\mathbb{P}}$
for $\mathbb{P}\in\mathcal{P}$ and $t\in[0,1]$.
(P1) $\Rightarrow$ (P2) in case each $\mathbb{P}$ is atomless: Let $A$ be an event with $\mathbb{P}(A)=1-\delta_{\mathbb{P}}$ such that $\{P>s\}\subseteq A\subseteq\{P\geq s\}$ for some $s\in[0,1]$. Note that $\mathbb{P}(P\leq t,A^{c})=\min\{\mathbb{P}(P\leq t),\mathbb{P}(A^{c})\}$. Therefore,
$\mathbb{P}(P\leq t,A)=(\mathbb{P}(P\leq t)-\mathbb{P}(A^{c}))_{+}\leq((1+\varepsilon_{\mathbb{P}})t+\delta_{\mathbb{P}}-\delta_{\mathbb{P}})_{+}=(1+\varepsilon_{\mathbb{P}})t$
for all $t\in(0,1)$, as desired.
(E2) $\Rightarrow$ (E1): It suffices to notice
$\mathbb{E}^{\mathbb{P}}[E\wedge t]=\mathbb{E}^{\mathbb{P}}[(E\wedge t)\mathds{1}_{A}]+\mathbb{E}^{\mathbb{P}}[(E\wedge t)\mathds{1}_{A^{c}}]\leq\mathbb{E}^{\mathbb{P}}[E\mathds{1}_{A}]+\mathbb{E}^{\mathbb{P}}[t\mathds{1}_{A^{c}}]\leq(1+\varepsilon_{\mathbb{P}})+\delta_{\mathbb{P}}t$
for $\mathbb{P}\in\mathcal{P}$ and $t\in\mathbb{R}_{+}$.
(E1) $\Rightarrow$ (E2) in case each $\mathbb{P}$ is atomless: Let $A$ be an event with $\mathbb{P}(A)=1-\delta_{\mathbb{P}}$ such that $\{E<t\}\subseteq A\subseteq\{E\leq t\}$ for some $t\in\mathbb{R}_{+}$. Note that $E\geq t$ on $A^{c}$. We have,
$\mathbb{E}^{\mathbb{P}}[E\mathds{1}_{A}]=\mathbb{E}^{\mathbb{P}}[(E\wedge t)\mathds{1}_{A}]=\mathbb{E}^{\mathbb{P}}[E\wedge t]-\mathbb{E}^{\mathbb{P}}[(E\wedge t)\mathds{1}_{A^{c}}]\leq 1+\varepsilon_{\mathbb{P}}+t\delta_{\mathbb{P}}-\mathbb{E}^{\mathbb{P}}[t\mathds{1}_{A^{c}}]\leq 1+\varepsilon_{\mathbb{P}},$
as desired. ∎
Without the atomless assumption on each $\mathbb{P}\in\mathcal{P}$, the implications (P1$\Rightarrow$(P2) and (E1$\Rightarrow$(E2) in Proposition 2.2 do not hold (one can find counterexamples on discrete spaces). We emphasize that the atomless assumption does not pertain to the distribution of $P$ or $E$ (i.e., $P$ or $E$ can be discrete). This atomless assumption can be avoided if external randomization is permitted.
The two formulations (P1) and (P2) in the above proposition are obviously equivalent when $\delta=0$ without requiring any condition; the same holds for (E1) and (E2).
Given an e-variable $E$, we can calibrate it into a p-variable via $P=(1/E)\land 1$, which is the only admissible choice of an e-to-p calibrator. Conversely, given a p-variable $P$, we can calibrate into an e-variable via a (p-to-e) calibrator in the sense of Vovk and Wang (2021). A (p-to-e) calibrator is a decreasing function $h:[0,\infty)\to[0,\infty]$ satisfying $\int_{0}^{1}h(x)\mathrm{d}x\leq 1$ and $h(x)=0$ for $x\in(1,\infty)$. Given such $h$ and a p-variable $P$, then $E=f(P)$ is an e-variable. The next theorem demonstrates that we can calibrate approximate p-variables and approximate e-variables in exactly the same way.
Theorem 2.3.
Let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$. Calibrating an $(\varepsilon,\delta)$-approximate p-variable yields an $(\varepsilon,\delta)$-approximate e-variable, and vice versa.
Proof..
Fix any $\mathbb{P}\in\mathcal{P}$. If $\delta_{\mathbb{P}}=1$ there is nothing to prove, so we assume $\delta_{\mathbb{P}}<1$ below. For the first statement, let $P$ be an $(\varepsilon,\delta)$-approximate p-variable for $\mathbb{P}$ and $f$ be a (p-to-e) calibrator and $b=(1+\varepsilon_{\mathbb{P}})/(1-\delta_{\mathbb{P}})\geq 1$. Let $F$ be a cdf on $[0,1/b]$ given by $F(s)=(1+\varepsilon_{\mathbb{P}})s+\delta_{\mathbb{P}}$ for $s\in[0,1/b]$. We have $\mathbb{P}(P\leq s)\leq F(s)$ for $s\in[0,1]$ and $F(0)=\delta_{\mathbb{P}}$. Hence, noting that the calibrator $f$ is decreasing, we have, for all $t\in\mathbb{R}_{+}$,
$\mathbb{E}^{\mathbb{P}}[f(P)\wedge t]$ $\leq\int_{0}^{1/b}(f(s)\wedge t)\mathrm{d}F(s)$
$\leq(1+\varepsilon_{\mathbb{P}})\int_{0}^{1/b}f(s)\mathrm{d}s+tF(0)$
$\leq(1+\varepsilon_{\mathbb{P}})\int_{0}^{1}f(s)\mathrm{d}s+\delta_{\mathbb{P}}t\leq 1+\varepsilon_{\mathbb{P}}+\delta_{\mathbb{P}}t.$
For the second statement, define $P=(1/E)\wedge 1$. We have, for $t\in(0,1)$,
$\mathbb{P}(P\leq t)$ $=\mathbb{P}(E\geq 1/t)$
$=\mathbb{P}(E\wedge(1/t)\geq 1/t)$
$\leq t\mathbb{E}^{\mathbb{P}}[E\wedge(1/t)]\leq t\left(1+\varepsilon_{\mathbb{P}}+\frac{\delta_{\mathbb{P}}}{t}\right)=(1+\varepsilon_{\mathbb{P}})t+\delta_{\mathbb{P}},$
as desired. ∎
Remark 2.4.
Following a similar proof, the statements in Theorem 2.3 hold true also for if we use the alternative formulations (P2) and (E2) in Proposition  2.2 for both approximate p-variables and approximate e-variables.
We end this section with the following comment. The roles of $(\varepsilon,\delta)$ are asymmetric and as mentioned earlier, the role of $\delta$ is more lenient. The following result shows that an $(\varepsilon,\delta)$-approximate p-variable (or e-variable) with $\varepsilon>0$ is always also an $(0,\delta^{\prime})$-approximate p-variable (or e-variable) where $\delta^{\prime}$ can be chosen strictly smaller than $\varepsilon+\delta$.
Proposition 2.5 ($\varepsilon$ to $\delta$).
Let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$, $\delta:\mathcal{P}\to[0,1]$, and $\delta^{\prime}=(\varepsilon+\delta)/(1+\varepsilon)$.
  1. (i)
    An ($\varepsilon,\delta$)-approximate p-variable for $\mathcal{P}$ is also an $(0,\delta^{\prime})$-approximate p-variable for $\mathcal{P}$.
  2. (ii)
    An ($\varepsilon,\delta$)-approximate e-variable for $\mathcal{P}$ is also an $(0,\delta^{\prime})$-approximate e-variable for $\mathcal{P}$.
Proof.
Let $P$ be an ($\varepsilon,\delta$)-approximate p-variable. Fix $\mathbb{P}\in\mathcal{P}$. Then, for any $t\in(0,(1-\delta_{\mathbb{P}})/(1+\varepsilon_{\mathbb{P}})),$ we have that:
$\mathbb{P}(P\leq t)\leq(1+\varepsilon_{\mathbb{P}})t+\delta\leq t+\varepsilon_{\mathbb{P}}\frac{1-\delta_{\mathbb{P}}}{1+\varepsilon_{\mathbb{P}}}+\delta_{\mathbb{P}}=t+\delta^{\prime}_{\mathbb{P}}.$
Meanwhile the inequality $\mathbb{P}(P\leq t)\leq t+\delta^{\prime}_{\mathbb{P}}$ is clearly true for $t\geq(1-\delta_{\mathbb{P}})/(1+\varepsilon_{\mathbb{P}})$ (since the right hand side is $\geq 1$). Thus $P$ is an $(0,\delta)^{\prime}$-approximate p-variable for $\mathcal{P}$.
The argument for the approximate e-variable is analogous (noting that only $t\geq 1$ needs to be checked in the condition for approximate e-variables) and omitted. ∎
Proposition 2.5 also suggests that controlling $\delta$ is a stronger requirement, leading to two different notions of asymptotic e-values and p-values in Section 3.
3 Asymptotic p-values and e-values
3.1 Definitions
In this section we consider a sequence of datasets $X=X^{(n)}$ of growing size as $n\to\infty$. Based on $X^{(n)}$ we compute a $[0,1]$-valued statistic $P^{(n)}=P^{(n)}(X^{(n)})$. Similarly, we can compute a nonnegative extended random variable $E^{(n)}=E^{(n)}(X^{(n)})$. The following definitions give some intuitive conditions on $P^{(n)}$ and $E^{(n)}$ for them to be considered as p-variables and e-variables in an asymptotic sense.
Definition 3.1 (Asymptotic p-variables and e-variables).
A sequence $(E^{(n)})_{n\in\mathbb{N}}$ of $[0,\infty]$-valued random variables is called a sequence of:
  • asymptotic e-variables for $\mathcal{P}$, if for each $n$, $E^{(n)}$ is $(\varepsilon_{n},\delta_{n})$-approximate for $\mathcal{P}$ for some $(\varepsilon_{n},\delta_{n})$, satisfying
    for each $\mbox{for each $\mathbb{P}\in\mathcal{P}$,}\quad\lim_{n\to\infty}\varepsilon_{n}(\mathbb{P})=\lim_{n\to\infty}\delta_{n}(\mathbb{P})=0;$ (1)
  • strongly asymptotic e-variables for $\mathcal{P}$, if condition (1) is replaced by
[…]
  • uniformly asymptotic e-variables for $\mathcal{P}$, if condition (1) is replaced by
    $\lim_{n\to\infty}\sup_{\mathbb{P}\in\mathcal{P}}\varepsilon_{n}(\mathbb{P})=\lim_{n\to\infty}\sup_{\mathbb{P}\in\mathcal{P}}\delta_{n}(\mathbb{P})=0;$
  • uniformly strongly asymptotic e-variables for $\mathcal{P}$, if condition (1) is replaced by
[…]
Sequences $(P^{(n)})_{n\in\mathbb{N}}$ of (uniformly, strongly) asymptotic p-variables for $\mathcal{P}$ are defined analogously.
We would also say “asymptotic e-values” (or p-values) in non-mathematical statements, and they can be interpreted as asymptotic e-variables (or p-variables) or their realizations. We sometimes simply say that a sequence is asymptotic for $\mathcal{P}$, and it should be clear from context whether we meant a sequence of asymptotic e-variables ($[0,\infty]$-valued) or that of asymptotic p-variables ($\mathbb{R}_{+}$-valued).
Clearly, if $\mathcal{P}$ is finite, then asymptotic e-variables or p-variables are also uniformly asymptotic.
3.2 Conditions for asymptotic p-values and e-values
We view the requirement of being strongly asymptotic as particularly intuitive for e-variables due the following result.
Proposition 3.2.
The sequence $(E^{(n)})_{n\in\mathbb{N}}$ of $[0,\infty]$-valued random variables is uniformly strongly asymptotic for $\mathcal{P}$ if and only if
$\limsup_{n\to\infty}\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}^{\mathbb{P}}[E^{(n)}]\leq 1.$
Similarly, it is strongly asymptotic for $\mathcal{P}$ if and only if $\limsup_{n\to\infty}\mathbb{E}^{\mathbb{P}}[E^{(n)}]\leq 1$ for all $\mathbb{P}\in\mathcal{P}$.
Proof.
We prove this result for the non-uniform case. The other argument is analogous. First suppose that $(E^{(n)})_{n\in\mathbb{N}}$ is a sequence of strongly asymptotic e-variables for $\mathcal{P}$. Fix $\mathbb{P}\in\mathcal{P}$. Then:
$\mathbb{E}^{\mathbb{P}}[E^{(n)}]\leq 1+\varepsilon_{n}(\mathbb{P})\to 1\text{ as }n\to\infty,$
where the first inequality holds for large enough $n$. For the other direction, note that any nonnegative extended random variable $E^{(n)}$ is an $(\varepsilon_{n},0)$-approximate e-variable for $\mathcal{P}$ with the choice $\varepsilon_{n}(\mathbb{P})=\max\{0,\,\mathbb{E}^{\mathbb{P}}[E^{(n)}]-1\}$ as long as $\mathbb{E}^{\mathbb{P}}[E^{(n)}]$ is finite. If $\limsup_{n\to\infty}\mathbb{E}^{\mathbb{P}}[E^{(n)}]\leq 1$, this implies that $\mathbb{E}^{\mathbb{P}}[E^{(n)}]<\infty$ for $n$ large enough and the above $\varepsilon_{n}(\mathbb{P})$ satisfies $\lim_{n\to\infty}\varepsilon_{n}(\mathbb{P})=0$, that is, $(E^{(n)})_{n\in\mathbb{N}}$ is a sequence of strongly asymptotic e-variables. ∎
The result below connects asymptotic e-variables and strongly asymptotic e-variables via uniform integrability.
Proposition 3.3.
If a sequence of asymptotic e-variables is uniformly integrable for each $\mathbb{P}\in\mathcal{P}$, then it is strongly asymptotic.
Proof.
Fix $\mathbb{P}\in\mathcal{P}$. Note that for $t\in\mathbb{R}_{+}$,
$\mathbb{E}^{\mathbb{P}}[E^{(n)}]\leq\mathbb{E}^{\mathbb{P}}[E^{(n)}\wedge t]+\mathbb{E}^{\mathbb{P}}[E^{(n)}\mathds{1}_{\{E^{(n)}\geq t\}}].$
By uniform integrability, for any $\eta>0$ there exists $T>0$ such that
$\sup_{n\in\mathbb{N}}\mathbb{E}^{\mathbb{P}}\left[E^{(n)}\mathds{1}_{\{E^{(n)}\geq T\}}\right]\leq\eta.$
Thus,
$\mathbb{E}^{\mathbb{P}}[E^{(n)}]\leq 1+\varepsilon_{n}(\mathbb{P})+\delta_{n}(\mathbb{P})T+\eta,$
and hence,
$\limsup_{n\to\infty}\mathbb{E}^{\mathbb{P}}[E^{(n)}]\leq 1+\eta.$
Since $\eta>0$ is arbitrary, we conclude by Proposition 3.2. ∎
The next proposition gives a sufficient condition that best describes the intuition behind asymptotic e-variables and p-variables.
Proposition 3.4.
If under each $\mathbb{P}\in\mathcal{P}$, a nonnegative sequence converges to an e-variable (resp. a p-variable) in distribution, then it is a sequence of asymptotic e-variables (resp. p-variables) for $\mathcal{P}$.
Proof.
Fix $\mathbb{P}\in\mathcal{P}$. Suppose that $(P^{(n)})_{n\in\mathbb{N}}$ converges in distribution to a p-variable $P$. Note that $\limsup_{n\to\infty}\mathbb{P}(P_{n}\leq t)\leq\mathbb{P}(P\leq t)\leq t$ for each $t\in[0,1]$. This implies $\lim_{n\to\infty}\mathbb{P}(P_{n}\leq t)\vee t=t$ for each $t\in[0,1]$. Since point-wise convergence of increasing functions to a continuous function on a compact interval implies uniform convergence, we have
$\lim_{n\to\infty}\sup_{t\in[0,1]}(\mathbb{P}(P_{n}\leq t)\vee t-t)=0\implies\lim_{n\to\infty}\sup_{t\in[0,1]}(\mathbb{P}(P_{n}\leq t)-t)_{+}=0.$
This shows that $(P^{(n)})_{n\in\mathbb{N}}$ is a sequence of asymptotic p-variables.
Next we show the statement on asymptotic e-variables. For $\delta\in[0,1]$, and a nonnegative random variable $E$, consider the function
$f_{\delta}(E)=\sup_{t\in\mathbb{R}_{+}}(\mathbb{E}^{\mathbb{P}}[E\wedge t]-\delta t)-1.$
By definition, if $\varepsilon:=f_{\delta}(E)\vee 0$ is finite, then $E$ is an $(\varepsilon,\delta)$-approximate e-variable for ${}\{\mathbb{P}\}$. Let $E$ be the limit of the nonnegative sequence $(E^{(n)})_{n\in\mathbb{N}}$ in distribution. Note that $f_{\delta}(E)\leq\mathbb{E}^{\mathbb{P}}[E]-1\leq 0$ for all $\delta\in[0,1]$. By the first-order condition, the supremum in $f_{\delta}(E)$ is attained for $t$ being a left or right $(1-\delta)$-quantile of $E$, and from there we can check that the function $f_{\delta}$ is continuous in convergence in distribution. This gives $\limsup_{n\to\infty}f_{\delta}(E^{(n)})\to f_{\delta}(E)\leq 0$. This implies, in particular, that for any $\delta>0$, there exists $N\in\mathbb{N}$ such that for $n>N$, $E^{(n)}$ is an $(\delta,\delta)$-approximate e-variable. Since $\mathbb{P}\in\mathcal{P}$ is arbitrary, this shows that $(E^{(n)})_{n\in\mathbb{N}}$ is a sequence of asymptotic e-variables for $\mathcal{P}$. ∎
Remark 3.5.
Gasparin et al. (2025) referred to asymptotic p-variables as those that converge to a p-variable in distribution. As Proposition 3.4 illustrates, this is a sufficient condition for our definition of asymptotic p-values. Our definition is more general. For instance, any sequence that is larger than a p-variable is an sequence of asymptotic p-variables, and it does not need to converge.
The next result gives equivalent conditions for uniformly asymptotic p-variables. The proof is similar to that of Proposition 3.4 and omitted here.
Proposition 3.6.
The sequence $(P^{(n)})_{n\in\mathbb{N}}$ of nonnegative random variables is uniformly asymptotic for $\mathcal{P}$ if and only if $\limsup_{n\to\infty}\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(P^{(n)}\leq t)\leq t$ for all $t\in(0,1)$. Similarly, it is asymptotic for $\mathcal{P}$ if and only if $\limsup_{n\to\infty}\mathbb{P}(P^{(n)}\leq t)\leq t$ for all $\mathbb{P}\in\mathcal{P}$ and $t\in(0,1)$.
We can naturally calibrate asymptotic p-variables and e-variables by applying a (possibly different) calibrator to each element in the sequence, in the sense of Theorem 2.3.
Proposition 3.7.
Calibrating a sequence of (uniformly, strongly) asymptotic p-variables yields a sequence of (uniformly, strongly) asymptotic e-variables, and vice versa.
This result follows directly from Theorem 2.3.
The most common way of constructing asymptotic p-variables and e-variables is via limit theorems such as the central limit theorem.
Proposition 3.8.
Let $Z^{(n)}$ be a statistic computed based on $X^{(n)}$ for $n\in\mathbb{N}$. Suppose that
$Z^{(n)}\stackrel{{\scriptstyle d}}{{\rightsquigarrow}}Z\,\text{ as }\,n\to\infty\,\text{ under every }\,\mathbb{P}\in\mathcal{P},$
where $Z$ is a random variable with a continuous cumulative distribution function $F$ and “$\stackrel{{\scriptstyle d}}{{\rightsquigarrow}}$” denotes convergence in distribution. Then the following hold:
  1. (i)
    Let $P^{(n)}=F(Z^{(n)})$ or $P^{(n)}=1-F(Z^{(n)})$ for $n\in\mathbb{N}$. Then $(P^{(n)})_{n\in\mathbb{N}}$ is a sequence of asymptotic p-variables for $\mathcal{P}$.
  2. (ii)
    Let $h:\mathbb{R}\to[0,\infty]$ be a function with $\mathbb{E}[h(Z)]\leq 1$ that is increasing, decreasing, or upper semi-continuous, and $E^{(n)}=h(Z^{(n)})$ for $n\in\mathbb{N}$. Then $(E^{(n)})_{n\in\mathbb{N}}$ is a sequence of asymptotic e-variables for $\mathcal{P}$. If $h$ is further bounded, then $(E^{(n)})_{n\in\mathbb{N}}$ is strongly asymptotic for $\mathcal{P}$.
Proof.
By the continuous mapping theorem, $P^{(n)}$ converges weakly to the uniform distribution, and hence item (i) directly follows from Proposition 3.4.
For (ii) we may argue as follows. If $h$ would have been continuous, then the result would follow from the continuous mapping theorem and Proposition 3.4. For upper semi-continuity, note that $\limsup_{n\to\infty}\mathbb{P}(E^{(n)}\geq x)\leq\mathbb{P}(h(Z)\geq x)$ for all $x\in\mathbb{R}$ by convergence in distribution. This gives $\limsup_{n\to\infty}f_{\delta}(E^{(n)})\leq f_{\delta}(h(Z))\leq 0$ using the notation and argument in the proof of Proposition 3.4, with some details omitted. The last statement follows from Proposition 3.3. When $h$ is increasing or decreasing, replacing it with its upper semi-continuous version, which does not change $\mathbb{E}[h(Z)]$, yields the desired statements. ∎
The claim for asymptotic p-variables of course is standard and underlies most constructions of p-values used in practice. But the statement for e-variables is new, which is our main motivation for stating the above result.
The most common cases in practice include the following textbook examples (Van Der Vaart, 1998; Lehmann and Romano, 2005):
  1. 1.
    $Z\sim\mathrm{N}(0,1)$ as would follow from the central limit theorem.
  2. 2.
    $Z\sim|\mathrm{N}(0,1)|$, where we use the notation $|\mathrm{N}(0,1)|$ to denote the half-normal distribution. This would follow also from the central limit theorem and the continuous mapping theorem. [^1] An application of this case is as follows. In the setting of Proposition 3.8(ii), suppose $Z\sim\mathrm{N}(0,1)$ but we seek to use a function $h$ that is increasing in $|z|$ (rather than $z$) and that satisfies $\mathbb{E}[h(|Z|)]\leq 1$. Then the sequence $(E^{(n)})_{n\in\mathbb{N}}$ with $E^{(n)}=h(|Z^{(n)}|)$ is also a sequence of asymptotic e-variables.
  3. 3.
    $Z\sim\chi^{2}_{\nu}$, where $\chi^{2}_{\nu}$ denotes the chi-square distribution with $\nu$ degrees of freedom. Such a limit occurs e.g., for likelihood ratio or chi-square tests.
3.3 A prototypical example of an asymptotic e-value
We next provide a prototypical example of asymptotic e-variables along the lines of the central limit theorem.
Example 3.9.
Suppose we observe the dataset $X^{(n)}=(X^{1},\dots,X^{n})$ where $X^{i}\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}$. We assume that their variance, $\mathrm{var}^{\mathbb{P}}(X^{i})$, is positive and finite, but is unknown. We seek to test the hypothesis
$\mathcal{P}=\{\mathbb{P}\in\mathcal{M}_{1}\,:\,\mathbb{E}^{\mathbb{P}}[X^{i}]=0,\,\mathrm{var}^{\mathbb{P}}({X^{i}})\in(0,\infty)\},$ (2)
where $\mathcal{M}_{1}$ denotes the set of all probability measures on $\mathbb{R}$. Define
$\bar{X}^{(n)}=\frac{1}{n}\sum_{i=1}^{n}X^{i},\;\;\;\,S^{(n)}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(X^{i})^{2}},\;\;\;\,\hat{\sigma}^{(n)}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}\left(X^{i}-\bar{X}^{(n)}\right)^{2}},$ (3)
and for fixed $\lambda\in\mathbb{R}$, let
$E^{(n)}=\exp\left(\lambda\frac{\sqrt{n}\bar{X}^{(n)}}{S^{(n)}}-\frac{\lambda^{2}}{2}\right),$ (4)
and
$\tilde{E}^{(n)}=\frac{1}{2}\exp\left(\lambda\frac{\sqrt{n}\bar{X}^{(n)}}{S^{(n)}}-\frac{\lambda^{2}}{2}\right)+\frac{1}{2}\exp\left(-\lambda\frac{\sqrt{n}\bar{X}^{(n)}}{S^{(n)}}-\frac{\lambda^{2}}{2}\right).$ (5)
The next proposition justifies that (4) defines a sequence of strongly asymptotic e-variables.
Theorem 3.10.
The sequences of random variables $E^{(n)}$ in (4) and $\tilde{E}^{(n)}$ in (5) are strongly asymptotic e-variables for $\mathcal{P}$ in (2). If $\hat{\sigma}^{(n)}$ is used in place of $S^{(n)}$ in (4) and (5), then $E^{(n)}$ and $\tilde{E}^{(n)}$ are asymptotic e-variables for $\mathcal{P}$. In either case, for $\lambda>0$, $E^{(n)}$ (resp. $\tilde{E}^{(n)}$) grows to infinity with probability $1$ under the alternative that $X^{1},\dots,X^{n}$ are drawn iid from $\mathbb{Q}$ with $\mathbb{E}^{\mathbb{Q}}[X^{i}]>0$ (resp. $\mathbb{E}^{\mathbb{Q}}[X^{i}]\neq 0$) and $\mathrm{var}^{\mathbb{Q}}(X^{i})<\infty$.
Proof..
Take any $\mathbb{P}\in\mathcal{P}$. By the central limit theorem, the law of large numbers and Slutsky’s theorem, $\sqrt{n}\bar{X}^{(n)}/S^{(n)}$ converges in distribution to $\mathrm{N}(0,1)$ as $n\to\infty$. We may verify that
$\mathbb{E}^{\mathbb{P}}\left[\exp\left(\lambda Z-\frac{\lambda^{2}}{2}\right)\right]=1,$
for $Z\sim\mathrm{N}(0,1)$. By Proposition 3.8 it follows that $E^{(n)}$ are asymptotic e-variables. To upgrade this result to strongly asymptotic, we can apply Theorem 2.5 of Giné et al. (1997) to show that the sequence $(E^{(n)})_{n\in\mathbb{N}}$ is uniformly integrable and then conclude the proof by Proposition 3.3.
Next take an alternative $\mathbb{Q}$ with $\mathbb{E}^{\mathbb{Q}}[X^{i}]>0$, $\mathrm{var}^{\mathbb{Q}}(X^{i})<\infty$, and suppose that $\lambda>0$. Observe the following. By the strong law of large numbers
$\bar{X}^{(n)}\to\mathbb{E}^{\mathbb{Q}}[X^{i}]>0\;\;\text{ almost surely as }\;\;n\to\infty,$
and
$S^{(n)}\to\sqrt{\mathrm{var}^{\mathbb{Q}}({X^{i}})+\mathbb{E}^{\mathbb{Q}}[X^{i}]^{2}}>0\;\;\text{ almost surely as }\;\;n\to\infty.$
The above imply that
$\lambda\frac{\sqrt{n}\bar{X}^{(n)}}{S^{(n)}}\to\infty\;\;\text{ almost surely as }\;\;n\to\infty,$
which also means that $E^{(n)}$ goes to $\infty$ with probability 1. The argument for $\tilde{E}^{(n)}$ is analogous and omitted. The argument for $\hat{\sigma}^{(n)}$ follows from Proposition 3.8. ∎
The above conclusion extends to a broader class of asymptotic e-variables obtained by mixing in (4) over $\lambda$ with respect to a distribution with bounded support. Concretely, let $G$ be any distribution on $\mathbb{R}$ with compact support. Then, under the assumptions of this section,
$\int\exp\left(\lambda\frac{\sqrt{n}\bar{X}^{(n)}}{S^{(n)}}-\frac{\lambda^{2}}{2}\right)\mathrm{d}G(\lambda),$
is a strongly asymptotic e-variable.
One central purpose of defining asymptotic e-variables is to handle cases where one does not know how to construct a nonasymptotic e-variable. Indeed, for the above $\mathcal{P}$, we do not know of nonasymptotic e-variables (other than constants). Had a bound on the variance been known, such nonasymptotic e-variables do exist, as shown next.
Example 3.11 (Testing the mean with bounded variance (Wang and Ramdas, 2023)).
Given some $\sigma>0$, consider the null hypothesis
$\mathcal{P}_{\sigma}=\left\{\mathbb{P}\in\mathcal{M}_{1}:\mathbb{E}^{\mathbb{P}}[X]=0,\;\mathrm{var}^{\mathbb{P}}({X})\leq\sigma^{2}\right\}.$
Then $X^{2}/\sigma^{2}$ is an e-variable for $\mathcal{P}_{\sigma}$, and for any $\lambda\in\mathbb{R}$ the following are e-variables for $\mathcal{P}_{\sigma}$:
$\exp\left(\lambda X-\frac{\lambda^{2}}{6}(X^{2}+2\sigma^{2})\right)$
and
$\exp\left(\phi(\lambda X)-\frac{\lambda^{2}}{2}\sigma^{2}\right),$
where $\phi(x)=\log(1+x+x^{2}/2)$ for $x\geq 0$ and $\phi(x)=-\log(1-x+x^{2}/2)$ for $x<0$. The different choices of e-variables have large e-power against different alternatives.
We leave the proof of the facts in Example 3.11 for the reader.
4 Compound p-values and e-values in multiple testing
We now extend the setting of Sections 2 and 3 to the multiple hypothesis testing setting. As before, we write $X$ to denote the entire dataset available to the researcher which is generated by some unknown probability distribution $\mathbb{P}^{*}$. We consider testing $K$ hypotheses determined by sets of probability measures $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$. For $k\in\mathcal{K}=\{1,\ldots,K\}$, the $k$-th hypothesis is defined as $H_{k}:\mathbb{P}^{*}\in\mathcal{P}_{k}$. We write
$\mathcal{N}(\mathbb{P})=\{k\in\mathcal{K}\,:\,\mathbb{P}\in\mathcal{P}_{k}\}\subseteq\mathcal{K},$
and hence $\mathcal{N}(\mathbb{P}^{*})$ denotes the indices of true null hypotheses.
We formally define the notion of compound p-variables and e-variables as fundamental statistical concepts. In what follows, $\mathcal{P}$ is a set of distributions that represents some assumptions or restrictions on the distributions in the testing problem.
Definition 4.1 (Compound p-variables and e-variables).
Fix the null hypotheses $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ and a set $\mathcal{P}$ of distributions.
  1. (i)
    Let $P_{1},\dots,P_{K}$ be nonnegative random variables. We say that $P_{1},\dots,P_{K}$ are compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$ if
    $\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{P}(P_{k}\leq t)\leq Kt\qquad\mbox{for all $t\in(0,1)$ and all $\mathbb{P}\in\mathcal{P}$.}$
  2. (ii)
    Let $E_{1},\dots,E_{K}$ be $[0,\infty]$-valued random variables. We say that $E_{1},\dots,E_{K}$ are compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$ if
    $\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[E_{k}]\leq K\qquad\mbox{for all $\mathbb{P}\in\mathcal{P}$.}$
    They are called tight compound e-variables if the supremum of the left hand side over $\mathbb{P}\in\mathcal{P}$ equals $K$.
We omit “under $\mathcal{P}$” in case $\mathcal{P}$ is the set of all distributions. Note that for all conditions to be checked in the above definition, it suffices to consider $\mathbb{P}\in\bigcup_{k\in\mathcal{K}}\mathcal{P}_{k}$. Therefore, it is without loss of generality to assume $\mathcal{P}\subseteq\bigcup_{k\in\mathcal{K}}\mathcal{P}_{k}$, and by default (without any restrictions) it is
$\mathcal{P}=\bigcup_{k\in\mathcal{K}}\mathcal{P}_{k}.$
Note that it is possible that $\mathbb{P}^{*}\notin\mathcal{P}$ as we do not know whether true null hypotheses exist.
When $K=1$, Definition 4.1 coincides with the standard definition of a single p-variable or e-variable. Of course, a vector of p-variables (resp. e-variables) is a trivial but important special case of compound p-variables (resp. e-variables). As is clear from the definition, no restriction is placed on the dependence structure of the p-variables or e-variables.
The notions in Definition 4.1 are not new. The notion of compound p-variables was introduced by Armstrong (2022) under a different name (average significance controlling p-values and tests) and further studied in an empirical Bayes setting by Ignatiadis and Sen (2025). Meanwhile, the notion of compound e-variables was already used in the literature as a technical tool, for instance, it was first used (without being given a name) in Wang and Ramdas (2022, Theorem 3). The term compound e-values first appeared in Ignatiadis, Wang, and Ramdas (2024) with a distribution-dependent upper bound of $|\mathcal{N}(\mathbb{P})|$ instead of $K$ (typically the value of $|\mathcal{N}(\mathbb{P})|$ is not known except $|\mathcal{N}(\mathbb{P})|\leq K$). Other authors have used the term generalized e-values (Banerjee et al., 2023; Bashari et al., 2023; Zhao and Sun, 2024; Lee and Ren, 2024) or relaxed e-values (Ren and Barber, 2024; Gablenz and Sabatti, 2024).
We can convert compound e-variables into compound p-variables and vice versa. If $E_{1},\dots,E_{K}$ are compound e-variables, then $(1/E_{1})\land 1,\dots,(1/E_{K})\land 1$ are compound p-variables, as can be seen by Markov’s inequality. Compound p-variables can be transformed into compound e-variables in the following result.
Theorem 4.2.
Let $h$ be any p-to-e calibrator. If $P_{1},\dots,P_{K}$ are compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$, then $E_{k}=h(P_{k}),\,k\in\mathcal{K},$ are compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$.
Proof.
Fix an atomless $\mathbb{P}$ and it suffices to verify the statement for this $\mathbb{P}$. Let $K_{0}=|\mathcal{N}(\mathbb{P})|$ and take $L\sim\mathrm{Unif}(\mathcal{N}(\mathbb{P}))$ and $U\sim\mathrm{Unif}[0,1]$, mutually independent and independent of everything else. Then
$\tilde{P}:=P_{L}\mathds{1}_{\{U\leq K_{0}/K\}}+\mathds{1}_{\{U>K_{0}/K\}}$
is a bona fide p-value because for $t\in(0,1)$,
$\mathbb{P}(\tilde{P}\leq t)=\mathbb{P}(P_{L}\leq t)\mathbb{P}(U\leq K_{0}/K)=\frac{K_{0}}{K}\frac{1}{K_{0}}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{P}(P_{k}\leq t)\leq t.$
Therefore,
$\frac{1}{K}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}\left[h(P_{k})\right]=\frac{K_{0}}{K}\mathbb{E}^{\mathbb{P}}\left[h(P_{L})\right]\leq\mathbb{E}^{\mathbb{P}}\left[h(P_{L})\mathds{1}_{\{U\leq K_{0}/K\}}+h(1)\mathds{1}_{\{U>K_{0}/K\}}\right]=\mathbb{E}^{\mathbb{P}}\left[h(\tilde{P})\right]\leq 1,$
where the last inequality holds since $h$ is a p-to-e calibrator. ∎
A first example of compound p-variables (resp. compound e-variables) is given by weighted p-variables (Genovese et al., 2006; Ramdas et al., 2019; Ignatiadis and Huber, 2021), resp. weighted e-variables (Wang and Ramdas, 2022).
Example 4.3 (Weighted p-variables and e-variables).
Let $w_{1},\ldots,w_{K}\geq 0$ be deterministic nonnegative numbers such that $\sum_{k\in\mathcal{K}}w_{k}\leq K$.
  1. (i)
    Let $P_{1}^{\prime},\dots,P_{K}^{\prime}$ be p-variables and define $\tilde{P}_{k}=P_{k}^{\prime}/w_{k}$ for $k\in\mathcal{K}$ with the convention $/0=0$ 0. Then, $\tilde{P}_{1},\dots,\tilde{P}_{K}$ are compound p-variables.
  2. (ii)
    Let $E_{1}^{\prime},\dots,E_{K}^{\prime}$ be e-variables and define $\tilde{E}_{k}={E}_{k}^{\prime}w_{k}$ for $k\in\mathcal{K}$ with the convention ×$\infty\times 0=\infty$. Then $\tilde{E}_{1},\dots,\tilde{E}_{K}$ are compound e-variables.
The class of compound p-variables (resp. e-variables) is much richer than the class of weighted p-variables (resp. e-variables) in Example 4.3, because the latter satisfy the stronger constraint
$\sum_{k\in\mathcal{K}}\sup_{\mathbb{P}\in\mathcal{P}_{k}}\mathbb{P}(\tilde{P}_{k}\leq t)\leq Kt\;\;\text{ for all }\,t\in(0,1),\,\,\qquad\text{resp.}\;\;\;\;\sum_{k\in\mathcal{K}}\sup_{\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[\tilde{E}_{k}]\leq K,$
compared to the condition for the compound p-variables/e-variables
$\sup_{\mathbb{P}\in\mathcal{P}}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{P}(P_{k}\leq t)\leq Kt\;\;\text{ for all }\,t\in(0,1),\,\,\qquad\text{resp.}\;\;\;\;\sup_{\mathbb{P}\in\mathcal{P}}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[E_{k}]\leq K.$
This follows from the fact that for any functions $x_{1},\dots,x_{K}:\mathcal{P}\to\mathbb{R}_{+}$,
$\sup_{\mathbb{P}\in\mathcal{P}}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}x_{k}(\mathbb{P})\leq\sum_{k\in\mathcal{K}}\sup_{\mathbb{P}\in\mathcal{P}_{k}}x_{k}(\mathbb{P}).$
We can combine p-variables and compound e-variables into compound p-variables as follows.
Example 4.4 (Combining p-variables and compound e-variables).
Suppose that $P_{k}$ is a p-variable for $\mathcal{P}_{k}$ for all $k\in\mathcal{K}$. $P_{1},\dots,P_{K}$ can be arbitrarily dependent. Moreover, let $E_{1},\dotsc,E_{K}$ be compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$. If $(E_{1},\dots,E_{K})$ is independent of $(P_{1},\dots,P_{K})$, then $P_{1}/E_{1},\dots,P_{K}/E_{K}$ are compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$. We revisit this example in Section 6.4.
We conclude this section with two general constructions of compound e-variables.
Example 4.5 (Convex combinations of compound e-variables).
Suppose that $E_{1}^{(\ell)},\dots,E_{K}^{(\ell)}$ are compound e-variables for $\ell=1,\ldots,L$. Let $w_{1},\dotsc,w_{L}\geq 0$ be deterministic nonnegative numbers such that $\sum_{\ell=1}^{L}w_{\ell}=1$. Define $E_{k}=\sum_{\ell=1}^{L}w_{\ell}E_{k}^{(\ell)}$. Then $E_{1},\dots,E_{K}$ are compound e-variables.
Example 4.6 (Combining evidence from follow-up studies).
Let $E_{1},\dots,E_{K}$ be compound e-variables calculated from data $X$. Based on these observations, suppose that we select a subset of indices $\mathcal{S}\subseteq\mathcal{K}$ of interest, for which we seek additional evidence. With this goal in mind, suppose we collect additional data $Y$ and summarize it in the form of e-variables $\{F_{j}\}_{j\in\mathcal{S}}$ that are conditionally valid given the past data; in particular, $\mathbb{E}^{\mathbb{P}}[F_{j}|E_{j}]\leq 1$ for all $j\in\mathcal{S}$ with $\mathbb{P}\in\mathcal{P}_{j}$. Then defining $F_{j}=1$ for $j\notin\mathcal{S}$, we have that $E_{1}F_{1},\dots,E_{K}F_{K}$ are also compound e-variables.
5 Approximate and asymptotic compound p/e-values
5.1 Approximate compound p/e-values
Just as in Definition 2.1 we introduced approximate p-variables and e-variables, below we define approximate compound p-variables and e-variables. The notion of approximate compound e-variables appears implicitly in Ignatiadis et al. (2024).
Definition 5.1 (Approximate compound p-variables and e-variables).
Fix the null hypotheses $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$, a set of distributions $\mathcal{P}$, and two functions $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$, and write $\varepsilon_{\mathbb{P}}=\varepsilon(\mathbb{P})$ and $\delta_{\mathbb{P}}=\delta(\mathbb{P})$ for $\mathbb{P}\in\mathcal{P}$.
  1. (i)
    Nonnegative random variables $P_{1},\dots,P_{K}$ are $(\varepsilon,\delta)$-approximate compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$ if
    $\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{P}(P_{K}\leq t)\leq K((1+\varepsilon_{\mathbb{P}})t+\delta_{\mathbb{P}})\qquad\mbox{for all $t\in(0,1)$and all $\mathbb{P}\in\mathcal{P}$.}$
  2. (ii)
    $[0,\infty]$-valued random variables $E_{1},\dots,E_{K}$ are $(\varepsilon,\delta)$-approximate compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$ if
    $\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[E_{k}\wedge t]\leq K(1+\varepsilon_{\mathbb{P}}+\delta_{\mathbb{P}}t)\qquad\mbox{for all $t\in\mathbb{R}_{+}$ and all $\mathbb{P}\in\mathcal{P}$.}$
If we start with per-hypothesis approximate p-variables (or e-variables), then we directly recover approximate compound p-variables (or e-variables) for the multiple testing problem.
Example 5.2.
Suppose that for all $k\in\mathcal{K}$, $P_{k}$ (resp. $E_{k}$) is an $(\varepsilon_{k},\delta_{k})$-approximate p-variable (resp. e-variable) for $\mathcal{P}_{k}$. Then, $P_{1},\dots,P_{K}$ (resp. $E_{1},\dots,E_{K}$) are $(\varepsilon,\delta)$-approximate compound p-variables (resp. e-variables) for the choice:
$\varepsilon=\frac{1}{K}\sum_{k\in\mathcal{K}}\varepsilon_{k},\qquad\delta=\frac{1}{K}\sum_{k\in\mathcal{K}}\delta_{k}.$
We can calibrate approximate compound p-variables to compound e-variables and vice versa as per usual if we use the same calibrator for all hypotheses. For instance, given a p-to-e calibrator $h$, if $P_{1},\dots,P_{K}$ are approximate compound p-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$ under $\mathcal{P}$, then $h(P_{1}),\dots,h(P_{k})$ are approximate compound e-variables for the same hypotheses.
Proposition 5.3.
Let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$. Calibrating $(\varepsilon,\delta)$-approximate compound p-variables yields $(\varepsilon,\delta)$-approximate compound e-variables, and vice versa.
Proof.
We only prove the calibration result from approximate compound p-variables to e-variables. To this end, let $P_{1},\dots,P_{K}$ be $(\varepsilon,\delta)$-approximate compound p-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$ and let $h$ be an e-to-p calibrator. Fix $\mathbb{P}\in\mathcal{P}$. Define $P_{L}$ and $\tilde{P}$ as in the proof of Theorem 4.2 and $\widetilde{\mathcal{P}}=\{\mathbb{P}\}$ as a singleton null hypothesis. Then, $\tilde{P}$ is an $(\varepsilon,\delta)$-approximate p-variable for $\widetilde{\mathcal{P}}$. Thus, by Theorem 2.3, $h(\tilde{P})$ is an $(\varepsilon,\delta)$-approximate e-variable for $\widetilde{\mathcal{P}}$.
Hence, further arguing as in the proof of Theorem 4.2, for any $t>0$,
$\frac{1}{K}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}\left[h(P_{k})\wedge t\right]=\frac{K_{0}}{K}\mathbb{E}^{\mathbb{P}}\left[h(P_{L}\wedge t)\right]\leq\mathbb{E}^{\mathbb{P}}\left[h(\tilde{P})\wedge t\right]\leq 1+\varepsilon_{\mathbb{P}}+\delta_{\mathbb{P}}t.$
Thus $h(P_{1}),\dots,h(P_{K})$ are $(\varepsilon,\delta)$-approximate e-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$. ∎
Just as Proposition 2.2 showed that approximate e-variables can be interpreted as allowing for a wiggle room in the definitions of p-variables and e-variables, a similar interpretation is also available for approximate compound p-variables and e-variables. We omit the proof as it is similar to the proof of Proposition 2.2.
Proposition 5.4.
Let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$ and $P_{1},\dots,P_{K}$ and $E_{1},\dots,E_{K}$ be nonnegative random variables ($[0,\infty]$-valued for $E_{k}$). For the following statements,
  1. (P1)
    $P_{1},\dots,P_{K}$ are $(\varepsilon,\delta)$-approximate compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$;
  2. (P2)
    for every $\mathbb{P}\in\mathcal{P}$, there exists an event $A$ such that
    $\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{P}(P_{k}\leq t,A)\leq K(1+\varepsilon_{\mathbb{P}})t~{}\mbox{for all $t\in(0,1)$ ~{}~{}and~{}~{} }\mathbb{P}(A)\geq 1-\delta_{\mathbb{P}};$
  3. (E1)
    $E_{1},\dots,E_{K}$ are $(\varepsilon,\delta)$-approximate compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$;
  4. (E2)
    for every $\mathbb{P}\in\mathcal{P}$, there exists an event $A$ such that
    $\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[E\mathds{1}_{A}]\leq K(1+\varepsilon_{\mathbb{P}})\mbox{ ~{}~{}and~{}~{} }\mathbb{P}(A)\geq 1-\delta_{\mathbb{P}};$
we have that (P2) implies (P1) and (E2) implies (E1). Moreover, if each $\mathbb{P}\in\mathcal{P}$ is atomless, then (P1)–(P2) are equivalent, and (E1)–(E2) are equivalent.
5.2 Asymptotic compound p/e-values
In this subsection we define asymptotic compound p-variables and e-variables. The reader may directly anticipate the definitions that follow. However, what is most important, is to clarify the asymptotic setup. As before, we seek to test hypotheses $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$ based on data $X^{(n)}$. We are interested in asymptotics in which either $n\to\infty$ (with $K$ fixed) or $K\to\infty$ (with $n$ fixed) or both $K,n\to\infty$.
Formally, we consider triangular-array type of asymptotics indexed by a single parameter $m\in\mathbb{N}$ that determines the dataset size, $n=n(m)$, and the number of hypotheses, $K=K(m)$. We require that $m\to\infty$ and that $\max\{K,n\}\to\infty$ as $m\to\infty$.
All three asymptotic regimes are frequently encountered in the (multiple) testing literature. The asymptotic regime with $n\to\infty$ and $K$ fixed is the setting we considered in Section 3 (with $K=1$), while Dudoit and van der Laan (2008) provide a textbook treatment of multiple testing in the regime with $n\to\infty$ and $K>1$ fixed. When using empirical Bayes arguments to share strength and learn nuisance parameters across hypotheses, one is relying on asymptotics as $K$ grows to infinity (as in e.g., the seminal FDR asymptotics in Storey et al. (2004)) and often with $n$ fixed (see Ignatiadis and Sen (2025) for a concrete example). Finally, other authors, e.g., Cao and Kosorok (2011); Liu and Shao (2014), have studied multiple testing procedures in the regime where both $n,K\to\infty$.
Definition 5.5 (Asymptotic compound p-variables and e-variables).
A sequence
$\left(E^{(n(m))}_{1},\dots,E^{(n(m))}_{K(m)}\right)_{m\in\mathbb{N}}$
of tuples of $[0,\infty]$-valued random variables is called a sequence of:
  • asymptotic compound e-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K(m)}$ under $\mathcal{P}$, if for each $m$, $E^{(n(m))}_{1},\dots,E^{(n(m))}_{K(m)}$ are $(\varepsilon_{m},\delta_{m})$-approximate compound e-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K(m)}$ under $\mathcal{P}$ for some $(\varepsilon_{m},\delta_{m})$, satisfying
    for each $\mbox{for each $\mathbb{P}\in\mathcal{P}$,}\quad\lim_{m\to\infty}\varepsilon_{m}(\mathbb{P})=\lim_{m\to\infty}\delta_{m}(\mathbb{P})=0;$ (6)
  • strongly asymptotic compound e-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K(m)}$ under $\mathcal{P}$, if condition (6) is replaced by
[…]
Uniformly (strongly) asymptotic compound e-variables are defined analogously. Moreover, sequences
$\left(P^{(n(m))}_{1},\dots,P^{(n(m))}_{K(m)}\right)_{m\in\mathbb{N}}$
of (uniformly, strongly) asymptotic compound p-variables are defined analogously.
5.3 The simultaneous t-test sequence model
We next describe a prototypical model of constructing asymptotic compound e-variables in a setting with composite null hypotheses. The construction appeared in Ignatiadis et al. (2024) and pertains to the problem of conducting multiple t-tests.
In our setting, we observe $K$ groups of data of size $n$, that is, we observe $X=(X_{k}^{1},\ldots,X_{k}^{n})_{k\in\mathcal{K}}$. We consider distributions $\mathbb{P}$ according to which all observations are jointly independent and the data within each group are iid. The universe of distributions $\mathcal{P}$ is understood to obey the above assumptions. Given any $\mathbb{P}\in\mathcal{P}$, we write $\mathbb{P}_{k}$ for its $k$-th group-wise marginal so that $X\sim\mathbb{P}$ entails that
$X_{k}^{i}\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}_{k}\;\;\text{ for }\;\;i=1,\dotsc,n.$
We seek to test the hypotheses,
$\mathcal{P}_{k}=\{\mathbb{P}\in\mathcal{P}\,:\,\mathbb{E}^{\mathbb{P}_{k}}[X_{k}^{i}]=0,\,\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})\in(0,\infty)\}.$
To this end, we start by constructing summary statistics in a group-wise manner, similar to (3):
$\bar{X}^{(n)}_{k}=\frac{1}{n}\sum_{i=1}^{n}X^{i}_{k},\;\;\;\,S^{(n)}_{k}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(X^{i}_{k})^{2}},\;\;\;\,\hat{\sigma}_{k}^{(n)}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}\left(X^{i}_{k}-\bar{X}_{k}^{(n)}\right)^{2}}.$ (7)
Ignatiadis et al. (2024) were interested in computing approximate compound e-variables that satisfy $E_{k}^{(n)}\propto(S_{k}^{(n)})^{2}$ for all $k\in\mathcal{K}$ [^2] The motivation was based on a normal sampling model with $X_{k}^{i}\,{\buildrel\text{iid}\over{\sim}\,}\mathrm{N}(\mu_{k},\sigma_{k}^{2})$. The natural test statistic for $H_{k}:\mu_{k}=0$ is the t-statistic $T_{k}=\sqrt{n}\bar{X}^{(n)}_{k}/{\hat{\sigma}^{(n)}_{k}}$, which by Basu’s theorem is independent of $(S_{k}^{(n)})^{2}$ under the null. The goal was to extract remaining information about $\mu_{k}$ through $(S_{k}^{(n)})^{2}$ before applying the ep-BH procedure (defined in Section 6.4 below) with the resulting asymptotic compound e-values and the t-test p-values. Such a procedure was shown to improve power compared to weighted multiple testing procedures using normalized weights. To motivate the construction that follows, fix any $\mathbb{P}\in\mathcal{P}$ and write $\mu_{k}(\mathbb{P})=\mathbb{E}^{\mathbb{P}_{k}}[X_{k}^{i}]$ and $\sigma_{k}^{2}(\mathbb{P})=\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})$ (supposing these exist). Observe that for any $k\in\mathcal{K}$,
$\mathbb{E}^{\mathbb{P}_{k}}[(S_{k}^{(n)})^{2}]=\sigma_{k}^{2}(\mathbb{P})+\mu_{k}^{2}(\mathbb{P})\,\text{ for all }\,k\in\mathcal{K}\quad\text{and so}\quad\mathbb{E}^{\mathbb{P}_{k}}[(S_{k}^{(n)})^{2}]=\sigma_{k}^{2}(\mathbb{P})\,\text{ for }\,\mathbb{P}\in\mathcal{P}_{k}.$
One can check that for any $\mathbb{P}\in(\bigcup_{k\in\mathcal{K}}\mathcal{P}_{k})\bigcap\mathcal{P}$,
$\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}\left[\tilde{E}_{k}^{(n)}(\mathbb{P})\right]=K,\quad\text{where}\quad\tilde{E}_{k}^{(n)}(\mathbb{P})=\frac{K\big{(}S_{k}^{(n)}\big{)}^{2}}{\sum_{j:\mathbb{P}\in\mathcal{P}_{j}}\sigma_{j}^{2}(\mathbb{P})}.$
The above construction is not directly usable in practice because of the dependence of $\tilde{E}_{k}^{(n)}(\mathbb{P})$ on $\mathbb{P}$. However, it motivates a construction in which we conservatively estimate the $\mathbb{P}$-dependent quantity $\sum_{j:\mathbb{P}\in\mathcal{P}_{j}}\sigma_{j}^{2}(\mathbb{P})$ by $\sum_{j\in\mathcal{K}}(\hat{\sigma}^{(n)}_{j})^{2}$ and then let
$E^{(n)}_{k}=\frac{K\big{(}S_{k}^{(n)}\big{)}^{2}}{\sum_{j\in\mathcal{K}}\big{(}\hat{\sigma}^{(n)}_{j}\big{)}^{2}},\quad k\in\mathcal{K}.$
We will show that the above are asymptotic compound e-variables under an asymptotic setup indexed by $m\in\mathbb{N}$ such that $n=n(m)$, $K=K(m)$ and $\max\{n,K\}\to\infty$ as $m\to\infty$. In case $K\to\infty$, we further restrict the class $\mathcal{P}$ to require that for any $\mathbb{P}\in\mathcal{P}$, |$|\{k\in\mathbb{N}:\mathbb{P}\in\mathcal{P}_{k}\}|=\infty$. In addition to this restriction (which remains implicit in our notation), our formal result will be stated in terms of the following two sets of distributions. The first set imposes some restrictions on moments of the marginals for null groups,
$\mathcal{P}^{\mathrm{B}}=\{\mathbb{P}\in\mathcal{P}\;:\;\mathbb{E}^{\mathbb{P}_{k}}[|X_{k}^{i}-\mathbb{E}^{\mathbb{P}_{k}}[X_{k}^{i}]|^{2(1+\delta)}]\leq C,\;\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})\geq c\;\text{ for all }\;k\;\text{ with }\;\mathbb{P}\in\mathcal{P}_{k}\},$
where $\delta,c,C\in(0,\infty)$ are fixed constants. The second set also imposes normality of the marginals for null groups,
$\mathcal{P}^{\mathrm{N}}=\{\mathbb{P}\in\mathcal{P}\;:\;\mathbb{P}_{k}=\mathrm{N}(0,\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})),\,\,c\leq\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})\leq C\;\text{ for all }\;k\;\text{ with }\;\mathbb{P}\in\mathcal{P}_{k}\}$
for some constants $c,C\in(0,\infty)$.
Proposition 5.6.
Suppose that $n=n(m)\geq 2$ and $K=K(m)\geq 1$ are such that $\max\{n,K\}\to\infty$ as $m\to\infty$. Then:
  1. (i)
    $(E_{1}^{(n)},\ldots,E_{K}^{(n)})_{m\in\mathbb{N}}$ is a sequence of asymptotic compound e-variables for $(\mathcal{P}_{1},\ldots,\mathcal{P}_{K})$ under $\mathcal{P}^{\mathrm{B}}$.
  2. (ii)
    $(E_{1}^{(n)},\ldots,E_{K}^{(n)})_{m\in\mathbb{N}}$ is a sequence of strongly asymptotic compound e-variables for $(\mathcal{P}_{1},\ldots,\mathcal{P}_{K})$ under $\mathcal{P}^{\mathrm{N}}$.
Proof.
Fix an arbitrary distribution $\mathbb{P}\in(\bigcup_{k}\mathcal{P}_{k})\bigcap\mathcal{P}^{\mathrm{B}}$ and write $\sigma_{k}^{2}=\sigma_{k}^{2}(\mathbb{P})$, omitting the dependence on $\mathbb{P}$. We write $K_{0}(\mathbb{P})=|\mathcal{N}(\mathbb{P})|=\{k\in\mathcal{K}:\mathbb{P}\in\mathcal{P}_{k}\}$. Note that $\max\{K_{0}(\mathbb{P}),n\}\to\infty$ and that
$\frac{1}{K}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}E_{k}^{(n)}\leq\frac{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}S_{k}^{(n)}\big{)}^{2}}{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}\hat{\sigma}^{(n)}_{k}\big{)}^{2}}.$ (8)
We will argue that the right hand side of the above expression converges in probability (and thus also in distribution) to the constant $1$. Arguing as in Proposition 3.4 it will then follow that $E_{1}^{(n)},\ldots,E_{K}^{(n)}$ are asymptotic compound e-variables. We claim that
$\frac{n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}S_{k}^{(n)}\big{)}^{2}}{n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\sigma_{k}^{2}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}1\;\;\text{ as }\;\;m\to\infty.$
Writing $U^{(m)}=n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}S_{k}^{(n)}\big{)}^{2}=\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\sum_{i=1}^{n}(X_{k}^{i})^{2}$, we see that $\mathbb{E}^{\mathbb{P}}[U^{(m)}]=n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\sigma_{k}^{2}$. Moreover, by our assumptions, there exists another constant $C^{\prime}$ that depends only on $\delta>0$ such that:
$\mathbb{E}^{\mathbb{P}}[U^{(m)}]\geq cnK_{0}(\mathbb{P}),\quad\mathbb{E}^{\mathbb{P}}[|U^{(m)}-\mathbb{E}^{\mathbb{P}}[U^{(m)}]|^{1+\delta}]\leq C^{\prime}CnK_{0}(\mathbb{P}).$
Thus, since $nK_{0}(\mathbb{P})\to\infty$ and using Markov’s inequality, it follows that $U^{(m)}/\mathbb{E}^{\mathbb{P}}[U^{(m)}]$ converges in probability to $1$ as desired. We can analogously argue that:
$\frac{n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}\hat{\sigma}^{(n)}_{k}\big{)}^{2}}{n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\sigma_{k}^{2}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}1\;\;\text{ as }\;\;m\to\infty.$
Thus implies that $E_{1}^{(n)},\ldots,E_{K}^{(n)}$ are asymptotic compound e-variables.
Next fix $\mathbb{P}\in(\bigcup_{k}\mathcal{P}_{k})\bigcap\mathcal{P}^{\mathrm{N}}$. It suffices to prove uniform integrability of the right hand side in (8) (using an argument analogous to the one of Proposition 3.3). We prove uniform integrability by controlling the following $\eta$-th moment for $\eta\in(1,1.25)$:
$\mathbb{E}^{\mathbb{P}}\left[\left(\frac{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}S_{k}^{(n)}\big{)}^{2}}{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}\hat{\sigma}^{(n)}_{k}\big{)}^{2}}\right)^{\eta}\right]\leq\mathbb{E}^{\mathbb{P}}\left[\left(\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}S_{k}^{(n)}\big{)}^{2}\right)^{2\eta}\right]^{1/2}\mathbb{E}^{\mathbb{P}}\left[\left(\frac{1}{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}\hat{\sigma}^{(n)}_{k}\big{)}^{2}}\right)^{2\eta}\right]^{1/2}.$
Notice that by our assumptions, $\sigma_{k}^{2}=\sigma_{k}^{2}(\mathbb{P})$ is both uniformly lower bounded and upper bounded for all $k$ such that $\mathbb{P}\in\mathcal{P}_{k}$. Ignoring multiplicative constants, it suffices to bound:
$\mathbb{E}^{\mathbb{P}}\left[\left(\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}n\big{(}S_{k}^{(n)}/\sigma_{k}\big{)}^{2}\right)^{2\eta}\right]^{1/2}\mathbb{E}^{\mathbb{P}}\left[\left(\frac{1}{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}(n-1)\big{(}\hat{\sigma}^{(n)}_{k}/\sigma_{k}\big{)}^{2}}\right)^{2\eta}\right]^{1/2}.$
This can be accomplished by noting that,
$\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}n\big{(}S_{k}^{(n)}/\sigma_{k}\big{)}^{2}\sim\chi^{2}_{nK_{0}(\mathbb{P})},\;\;\;\;\;\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}(n-1)\big{(}\hat{\sigma}^{(n)}_{k}/\sigma_{k}\big{)}^{2}\sim\chi^{2}_{(n-1)K_{0}(\mathbb{P})},$
where $\chi^{2}_{d}$ denotes the chi-square distribution with $d$ degrees of freedom. The remainder of the calculation amounts to direct evaluation of the $2\eta$-th moment (resp. negative $2\eta$-th moment) of chi squared random variables. ∎
6 Compound e-values are central to multiple testing
Our goal in this section is to demonstrate that compound e-values are fundamental to multiple testing and provide a useful conceptual framework for developing new multiple testing procedures that control the false discovery rate. To motivate what follows, we first record an observation about single hypothesis testing in which we seek to construct a level-$\alpha$ non-randomized test for a hypothesis $\mathcal{P}$, that is, a $\{0,1\}$-valued function $\phi(\cdot)$ such that $\mathbb{E}^{\mathbb{P}}[\phi(X)]\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}$. First, if $E=E(X)$ is an e-value for $\mathcal{P}$, then $\phi(X)=\mathds{1}_{\{E(X)\geq 1/\alpha\}}$ is a level-$\alpha$ test, since by Markov’s inequality $\mathbb{P}(E(X)\geq 1/\alpha)\leq\alpha\mathbb{E}^{\mathbb{P}}[E(X)]\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}$. Moreover, given any level-$\alpha$ test $\phi(X)$, we can construct an e-value via $E=\phi(X)/\alpha$ in which case $E\geq 1/\alpha$ if and only if $\phi(X)=1$. Thus, every test (including those based on thresholding p-values) implicitly uses an e-value and is recoverable by thresholding e-values at $1/\alpha$.
The remainder of this section develops a generalization of the above relation in the context of multiple testing. In Section 6.1 we show that compound e-values underlie every procedure controlling the false discovery rate (FDR), while in Section 6.2 we show that every FDR procedure can be recovered by applying the e-BH procedure of Wang and Ramdas (2022) to compound e-values. We then continue with practical applications of compound e-values to combination and derandomization of multiple testing procedures (Section 6.3) and explain why compound e-values provide the natural notion of weights for the weighted p-BH procedure (Section 6.4). We conclude by presenting a few known results on multiple testing with compound p-values (Section 6.5).
6.1 Compound e-values underlie every FDR procedure
We first introduce some further notation. We denote a multiple testing procedure by $\mathcal{D}$, that is, a Borel function of $X$ that produces a subset of $\mathcal{K}$ representing the indices of rejected hypotheses. The rejected hypotheses by $\mathcal{D}$ are called discoveries. We write $V_{k}=\mathds{1}_{\{k\in\mathcal{D}\}}$ as the indicator of whether $\mathcal{P}_{k}$ is rejected by $\mathcal{D}$, $F_{\mathcal{D}}=F_{\mathcal{D}}(\mathbb{P})$ as the number of true null hypotheses that are rejected (i.e., false discoveries) when the true distribution is $\mathbb{P}$:
$F_{\mathcal{D}}=\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}V_{k},$
and $R_{\mathcal{D}}$ as the total number of discoveries:
$R_{\mathcal{D}}=\sum_{k\in\mathcal{K}}V_{k}.$
Benjamini and Hochberg (1995) proposed to control the false discovery rate, which is the expected value of the false discovery proportion, that is, $\mathrm{FDR}_{\mathcal{D}}^{\mathbb{P}}:=\mathbb{E}^{\mathbb{P}}[{F_{\mathcal{D}}}/{R_{\mathcal{D}}}]$ with the convention $/0=0$ 0.
Throughout this section, we assume that each $\mathbb{P}$ is atomless; this means that we can simulate a uniformly distributed random variable under $\mathbb{P}$. This allows us to use the equivalent definition in Proposition 2.2 of approximate e-variables through an event $A$ with a constraint on its probability.
Definition 6.1 (Exact, approximate, and asymptotic FDR control).
Fix the null hypotheses $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ and a set $\mathcal{P}$ of distributions. Let $\mathcal{D}$ be a multiple testing procedure and let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$. The procedure $\mathcal{D}$ is said to have
  • FDR control at level $\alpha$ for $(\mathcal{P}_{1},\dotsc,\mathcal{P}_{K})$ under $\mathcal{P}$ if for every $\mathbb{P}\in\mathcal{P}$
    $\mathrm{FDR}_{\mathcal{D}}^{\mathbb{P}}=\mathbb{E}^{\mathbb{P}}\left[\frac{F_{\mathcal{D}}}{R_{\mathcal{D}}}\right]\leq\alpha;$
  • $(\varepsilon,\delta)$-approximate FDR control at level $\alpha$ for $(\mathcal{P}_{1},\dotsc,\mathcal{P}_{K})$ under $\mathcal{P}$, if for every $\mathbb{P}\in\mathcal{P}$, there exists an event $A$ such that
    $\mathbb{E}^{\mathbb{P}}\left[\frac{F_{\mathcal{D}}}{R_{\mathcal{D}}}\mathds{1}_{A}\right]\leq\alpha(1+\varepsilon_{\mathbb{P}})~{}\mbox{~{}~{}and~{}~{} }\mathbb{P}(A)\geq 1-\delta_{\mathbb{P}};$
  • asymptotic FDR control at level $\alpha$ for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K(m)})$ under $\mathcal{P}$ if for each $m\in\mathbb{N}$ (with $m$ indexing a sequence of problems as in Section 5.2) it has $(\varepsilon_{m},\delta_{m})$-approximate FDR control at level $\alpha$ for some $(\varepsilon_{m},\delta_{m})$, satisfying
    for each $\mbox{for each $\mathbb{P}\in\mathcal{P}$,}\quad\lim_{m\to\infty}\varepsilon_{m}(\mathbb{P})=\lim_{m\to\infty}\delta_{m}(\mathbb{P})=0.$
We make a few remarks. A notion related to approximate FDR control was used as proof technique (without being given that name) by Blanchard and Roquain (2008). A procedure with $(\varepsilon,\delta)$-approximate FDR control also satisfies
$\mathrm{FDR}_{\mathcal{D}}^{\mathbb{P}}\leq\alpha(1+\varepsilon_{\mathbb{P}})+\delta_{\mathbb{P}}\text{ for any }\mathbb{P}\in\mathcal{P}.$
Conversely, any procedure satisfying the above inequality has $(\varepsilon,\delta/\alpha)$-approximate FDR control, by invoking the equivalence in Proposition 2.2; more precisely, the above condition implies, for all $t\geq 1$,
$\mathbb{E}^{\mathbb{P}}\left[\frac{F_{\mathcal{D}}}{\alpha{R_{\mathcal{D}}}}\wedge t\right]\leq\mathbb{E}^{\mathbb{P}}\left[\frac{F_{\mathcal{D}}}{\alpha{R_{\mathcal{D}}}}\right]\leq(1+\varepsilon_{\mathbb{P}})+\frac{\delta_{p}}{\alpha}\leq(1+\varepsilon_{\mathbb{P}})+\frac{\delta_{\mathbb{P}}}{\alpha}t.$
The following result enables the construction of (approximate/asymptotic) compound e-values from any (approximate/asymptotic) FDR controlling procedure. The key observation underlying it has been made by several authors (Ren and Barber, 2024; Banerjee et al., 2023; Li and Zhang, 2023, 2025) for exact nonasymptotic FDR control.
Theorem 6.2 (From FDR control to compound e-values).
Let $\mathcal{D}$ be any procedure that controls the false discovery rate at a known level $\alpha\in(0,1]$ for $(\mathcal{P}_{1},\dotsc,\mathcal{P}_{K})$ under $\mathcal{P}$. Define
$E_{k}=\frac{K}{\alpha}\frac{V_{k}}{R_{\mathcal{D}}\lor 1}.$ (9)
Then, $E_{1},\dots,E_{K}$ are compound e-variables for $(\mathcal{P}_{1},\dotsc,\mathcal{P}_{K})$ under $\mathcal{P}$. Further, if FDR control only holds $(\varepsilon,\delta)$-approximately or asymptotically, then $E_{1},\dots,E_{K}$ are $(\varepsilon,\delta)$-approximate or asymptotic compound e-values.
Proof.
Fix any $\mathbb{P}\in\mathcal{P}$. The result can then be seen from
$\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[E_{k}]=\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}\left[\frac{K}{\alpha}\frac{V_{k}}{R_{\mathcal{D}}\lor 1}\right]=\frac{K}{\alpha}\mathrm{FDR}_{\mathcal{D}}^{\mathbb{P}}\leq\frac{K}{\alpha}\alpha=K,$
where the inequality follows from the definition of FDR guarantee. The proof of the approximate and asymptotic claims follows analogously. ∎
6.2 Every FDR procedure is an instance of e-BH
Let $E_{1},\dots,E_{K}$ be compound e-values. Then we can control the FDR through the e-BH procedure, which is defined as follows.
Definition 6.3 (e-BH procedure, Wang and Ramdas, 2022).
Let $E_{1},\dots,E_{K}$ be $[0,\infty]$-valued random variables. For $k\in\mathcal{K}$, let $E_{[k]}$ be the $k$-th order statistic of $E_{1},\ldots,E_{K}$, from the largest to the smallest. The e-BH procedure rejects all hypotheses with the largest $k_{e}^{*}$ e-values, where
$k_{e}^{*}:=\max\left\{k\in\mathcal{K}:\frac{kE_{[k]}}{K}\geq\frac{1}{\alpha}\right\},$
with the convention $\max(\varnothing)=0$.
The following theorem due to Wang and Ramdas (2022) shows that e-BH controls the FDR.
Theorem 6.4.
Suppose that we apply the e-BH procedure at level $\alpha$ to $E_{1},\dots,E_{K}$. Then:
  • If $E_{1},\dots,E_{K}$ are compound e-values, then e-BH controls the false discovery rate at level $\alpha$.
  • If $E_{1},\dots,E_{K}$ are $(\varepsilon,\delta)$-approximate compound e-values, then e-BH satisfies $(\varepsilon,\delta)$-approximate FDR control, and so it controls the FDR at level $\alpha(1+\varepsilon)+\delta$.
  • If $E_{1},\dots,E_{K}$ are asymptotic compound e-values, then e-BH has asymptotic FDR control at level $\alpha$.
Notably, the above result requires no assumption whatsoever on the dependence across the $E_{k}$. The results in Wang and Ramdas (2022) handled $(\varepsilon,0)$-approximate compound e-values, so technically Theorem 6.4 offers a slight generalization in that respect.
Several authors including Ren and Barber (2024); Li and Zhang (2023, 2025) have observed that specific multiple testing procedures may be recast as e-BH with a specific choice of compound e-values as constructed in Theorem 6.2. In fact one can show, as we do next, that this holds for any multiple testing procedure that controls the false discovery rate at a known level. This implies that if one seeks to control the FDR, it is without loss of generality to search among all procedures that apply e-BH to compound e-values.
In what follows, an FDR procedure $\mathcal{D}$ at level $\alpha$ is admissible if for any FDR procedure $\mathcal{D}^{\prime}$ at level $\alpha$ such that $\mathcal{D}\subseteq\mathcal{D}^{\prime}$, we have $\mathcal{D}=\mathcal{D}^{\prime}$.
Theorem 6.5 (Universality of e-BH with compound e-values).
Let $\mathcal{D}$ be any procedure that controls the FDR at a known level $\alpha$ for the hypotheses $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$ under $\mathcal{P}$. Then, there exists a choice of compound e-values such that the e-BH procedure yields identical discoveries as $\mathcal{D}$. Further, if the original FDR procedure is admissible, then these compound e-values can be chosen to be tight.
Proof.
The first part directly follows from Theorems 6.2 and 6.4. To show the last statement, consider an FDR procedure $\mathcal{D}$ at level $\alpha$ and take the compound e-values $E_{1},\dots,E_{K}$ from (9). Define
$K^{*}:=\sup_{\mathbb{P}\in\mathcal{P}}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[E_{k}].$
If $K^{*}$ is equal to $K$, then there is nothing to show as $E_{1},\dots,E_{K}$ are tight compound e-values. Otherwise, $K^{*}<K$. The case $K^{*}=0$ means one never rejects any hypotheses, for which choosing $E_{1}^{\prime}=\dots=E_{K}^{\prime}=1$ would suffice as tight compound e-values that produce $\mathcal{D}$ via e-BH. For $K^{*}>0$, we let
$E_{k}^{\prime}=\frac{K}{K^{*}}E_{k}\mbox{~{}~{}~{}for $k\in\mathcal{K}$},$
and apply the e-BH procedure to $E^{\prime}_{1},\dots,E^{\prime}_{K}$ (which are tight compound e-values). This new procedure controls FDR at level $\alpha$ and produces at least as many discoveries as $\mathcal{D}$. Since $\mathcal{D}$ is admissible, this procedure must coincide with $\mathcal{D}$, and hence $\mathcal{D}$ is e-BH applied to the tight compound e-values $E^{\prime}_{1},\dots,E^{\prime}_{K}$. ∎
6.3 Combination and derandomization
The connections between FDR controlling procedures, e-BH, and compound e-values laid out in Sections 6.1 and 6.2 motivate the following general and practical mechanism for combining discoveries across multiple testing procedures. To be concrete, suppose we run $L$ different multiple testing procedures $\mathcal{D}_{1},\dots,\mathcal{D}_{L}$. The $\ell$-th procedure is applied to a subset of hypotheses $\{\mathcal{P}_{k}:\;k\in\mathcal{\mathcal{missing}}{S}_{\ell}\}$ where $\mathcal{S}_{\ell}\subseteq\mathcal{K}$ and controls the FDR at level $\alpha_{\ell}$. Then we may proceed as follows:
  1. 1.
    For the $\ell$-th multiple testing procedure, form compound e-values $E_{k}^{(\ell)},\,k\in\mathcal{S}_{\ell}$, for $\mathcal{P}_{k},\,k\in\mathcal{S}_{\ell}$ (which always exist by Theorem 6.2). They could be—but need not necessarily be—the implied ones formed in Theorem 6.2. For $k\in\mathcal{K}\setminus\mathcal{S}_{\ell}$, set $E_{k}^{(\ell)}=1$.
  2. 2.
    Fix weights $w_{1},\dots,w_{L}\geq 0$ with $\sum_{\ell=1}^{L}w_{\ell}=1$. Then construct new compound e-values $E_{1},\dots,E_{K}$ by convex combination as in Example 4.5, i.e., $E_{k}=\sum_{\ell=1}^{L}w_{\ell}E_{k}^{(\ell)}$ for all $k\in\mathcal{K}$. We also allow $w_{1},\dots,w_{L}\geq 0$ to be random, as long as they are independent of all e-values used in the procedures, and in that case it suffice to require $\sum_{\ell=1}^{L}\mathbb{E}^{\mathbb{P}}[w_{\ell}]\leq 1$ for all $\mathbb{P}\in\mathcal{P}$ (this condition is similar to the condition for compound e-values).
  3. 3.
    Apply the e-BH procedure to the new compound e-values $E_{1},\dots,E_{K}$ at level $\alpha$.
The above construction is guaranteed to control the FDR at level $\alpha$, as long as all the individual procedures $\mathcal{D}_{\ell}$ form valid compound e-values. Note that the value of $\alpha_{\ell}$ does not matter, as it is only used in the construction of $E_{1}^{(\ell)},\dots,E_{K}^{(\ell)}$ , possibly implicitly (e.g., in Theorem 6.2).
One important application of the above recipe is derandomization. Suppose that $\mathcal{D}_{\ell}$ is a randomized multiple testing procedure, that is, it is a function of both the data $X$ as well as a random variable $U_{\ell}$ generated during the analysis. Such randomness may not be desirable, since different random number generation seeds will lead to different sets of discoveries. In such cases, the above recipe can be used to construct a new derandomized procedure $\mathcal{D}$ that is less sensitive to $U_{1},\dots,U_{L}$ (formally, full derandomization would occur if $U_{\ell}$ are iid and $L\to\infty$). In this way Ren and Barber (2024) were able to derandomize the model-X knockoff filter (Candès et al., 2018), a flexible set of methods for variable selection in regression with finite-sample FDR control that previously relied on additional randomness.
Some further applications include the following: Banerjee et al. (2023) used the above recipe for meta-analysis in which the $\ell$-th study only reports the set of tested hypotheses, the set of discoveries, as well as the targeted FDR level. Li and Zhang (2023, 2025) considered several applications including the development of new methods for multiple testing with side information.
6.4 Compound e-values as weights in p-value based multiple testing
Ignatiadis et al. (2024) explained that e-values can be used as weights in p-value based multiple testing. We explain why compound e-values are the natural notion of weights for multiple testing with the p-BH procedure. We first define the p-BH procedure of Benjamini and Hochberg (1995) for FDR control based on p-values, as well as weighted generalizations thereof.
Definition 6.6 (p-BH, weighted p-BH, ep-BH).
Let $P_{1},\dots,P_{K}$ be p-values for the hypotheses $H_{1},\dots,H_{K}$ and let $W_{1},\dots,W_{K}$ be $[0,\infty]$-valued. Define $Q_{k}=P_{k}/W_{k}$ with the convention $/0=0$ 0. For $k\in\mathcal{K}$, let $Q_{(k)}$ be the $k$-th order statistic of $Q_{1},\ldots,Q_{K}$, from the smallest to the largest. Consider the procedure that rejects hypothesis $k$ if $Q_{k}$ is among the smallest $k_{q}^{*}$ values $Q_{1},\dots,Q_{K}$, where
$k_{q}^{*}:=\max\left\{k\in\mathcal{K}:\frac{KQ_{(k)}}{k}\leq\alpha\right\},$
with the convention $\max(\varnothing)=0$. When $W_{k}=1$ for all $k$, then the above procedure is the p-BH procedure of Benjamini and Hochberg (1995). When $W_{k}=w_{k}$ is deterministic and such that $\sum_{k\in\mathcal{K}}w_{k}=K$, then the above procedure is the weighted p-BH procedure of Genovese et al. (2006). When $W_{k}=E_{k}$ and $E_{1},\dots,E_{K}$ are compound e-values, then the above procedure is the ep-BH (e-weighted p-BH) procedure of Ignatiadis et al. (2024).
The generalization of p-BH to allow for deterministic weights by Genovese et al. (2006) was motivated by differentially prioritizing hypotheses to improve the power of p-BH while maintaining frequentist FDR control. The requirement $\sum_{k\in\mathcal{K}}w_{k}=K$ may be interpreted as a fixed size budget to be split across hypotheses. When the p-values are positive regression dependent on a subset (PRDS, Benjamini and Yekutieli (2001)) under the distributions in $\mathcal{P}$ and the weights are deterministic then existing arguments in the literature (Ramdas et al., 2019) demonstrate that the FDR of the weighted p-BH procedure is controlled at $(\alpha/K)\sum_{k\in\mathcal{N}}w_{k}$ under this assumption, which is bounded by $\alpha$ under the budget constraint $\sum_{k\in\mathcal{K}}w_{k}=K$. If the weights $W_{1},\dots,W_{K}$ are random and independent of $P_{1},\dots,P_{K}$, then by applying the above argument conditional on the weights and by iterated expectation, it follows that the FDR is controlled at $\sup_{\mathbb{P}\in\mathcal{P}}(\alpha/K)\mathbb{\sum}_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[W_{k}]$. The latter is controlled at $\alpha$ precisely when $W_{1},\dots,W_{K}$ are compound e-values. We note that Ignatiadis et al. (2024) defined the ep-BH procedure and proved FDR control (in their Theorem 4) when $E_{1},\dots,E_{K}$ are e-values. However, in Section 4.5, Ignatiadis et al. (2024) explained that the guarantees would continue to hold for compound e-values (although their definition is slightly different from ours).
In Appendix A we also show that compound e-variables can be used as weights for merging p-variables.
6.5 Compound p-values in multiple testing
So far in this section we have focused on the role of compound e-values in multiple testing. Here we briefly discuss some known results about compound p-values. Armstrong (2022) provided a counterexample of jointly independent compound p-values $P_{1},\dots,P_{K}$, such that the p-BH procedure (Definition 6.6) does not control the false discovery rate.
On the other hand, we show the following result, slightly generalizing a claim by Armstrong (2022).
Proposition 6.7.
If $P_{1},\dots,P_{K}$ are (asymptotic) compound p-values, then the p-BY (Benjamini and Yekutieli, 2001) procedure (that is, the p-BH procedure applied at level $\alpha^{\prime}=\alpha\sum_{k\in\mathcal{K}}1/k$) provides (asymptotic) control of the false discovery rate at $\alpha$ under arbitrary dependence among $P_{1},\dots,P_{K}$.
Proof.
The nonasymptotic result was proved by Armstrong (2022), but here is an alternate proof that makes it transparent to extend to the asymptotic case. Wang and Ramdas (2022) proved that the p-BY procedure may be equivalently described by two steps. First, the p-values $P_{k}$ are calibrated to e-values $E_{k}$ via a specific choice of calibrator $f:p\mapsto T(\alpha/(\ell_{K}p))/\alpha$, where
$T(x)=\frac{K}{\lceil K/x\rceil}\mathds{1}_{\{x\geq 1\}}\mbox{~{}~{}with $T(\infty)=K$}.$
Second, the e-BH procedure is applied to $E_{1},\dots,E_{K}$. With this, the nonasymptotic claim then follows directly from Theorems 4.2 and 6.4. The asymptotic claim follows because calibrating asymptotic p-values yields asymptotic e-values, which when used with the e-BH procedure provides asymptotic FDR control. ∎
Furthermore, Armstrong (2022) proved that the p-BH procedure with compound p-values controls the false discovery rate under weak dependence asymptotics (as in Storey et al. (2004)); such a property of the p-BH procedure (with compound p-values) was also anticipated earlier by Efron (2007).
7 Sequence models and compound decisions
In this section we focus on sequence models, wherein the data $X$ may be written as $X=(X_{k}:k\in\mathcal{K})$ and in principle we may test each hypothesis $H_{k}$ using only $X_{k}$. We record the following definition, which we will call upon throughout this section.
Definition 7.1 (Simple separable e-values).
$E_{1},\dots,E_{K}$ are called separable if for all $k$, $E_{k}$ is $X_{k}$-measurable, that is $E_{k}=E_{k}(X_{k})$. They are called simple separable if $E_{k}=f(X_{k})$ for some function $f$ (which is the same for all $k$).
In what follows, we provide a brief summary of compound decision theory which motivates the nomenclature “compound e-values” (Section 7.1) and then we provide two constructions of (approximate) compound e-values motivated by compound decision theory (Sections 7.2 and 5.3).
7.1 Compound decision theory
Why do we use the term compound e-values? Our motivation is to pay tribute and connect the definition to Robbins’ compound decision theory (Robbins, 1951) in which multiple statistical problems are connected through a (compound) loss function that averages over individual losses. We refer the reader to e.g. Copas (1969); Zhang (2003); Jiang and Zhang (2009) for more comprehensive accounts and present a brief overview here. As a very concrete example (which was studied in detail already in Robbins (1951)), consider the Gaussian sequence model:
$\boldsymbol{\mu}=(\mu_{1},\dotsc,\mu_{K})\text{ fixed},\quad X_{k}\,{\buildrel\text{ind}\over{\sim}\,}\mathrm{N}(\mu_{k},1)\text{ for }k\in\mathcal{K}.$ (10)
Robbins was interested in constructing estimators $\hat{\mu}_{k}$ of $\mu_{k}$ such that the following expected compound loss would be small:
$\frac{1}{K}\sum_{k\in\mathcal{K}}\mathbb{E}[(\hat{\mu}_{k}-\mu_{k})^{2}].$ (11)
If $\mu_{k}\,{\buildrel\text{iid}\over{\sim}\,}M$ for some prior distribution $M$, then the best estimator is the Bayes estimator $\hat{\mu}_{k}^{B}=\mathbb{E}^{M}{[\mu_{k}\mid X]}=\mathbb{E}^{M}{[\mu_{k}\mid X_{k}]}$, which is simple separable (in the sense of Definition 7.1). [^3] The conditional expectations in the above displays refer to the joint distribution of $(\mu_{k},X_{k}),\,k\in\mathcal{K}$. Robbins asked: Suppose $\boldsymbol{\mu}$ is deterministic as in (10). Given knowledge of $\boldsymbol{\mu}$, which function $s_{\boldsymbol{\mu}}:\mathbb{R}\to\mathbb{R}$ leads to a simple separable estimator $\hat{\mu}_{k}^{s_{\boldsymbol{\mu}}}=s_{\boldsymbol{\mu}}(X_{k})$ that minimizes the risk in (11)? The answer lies in the fundamental theorem of compound decisions, which formally connects (10) to the univariate Bayesian problem with prior $M(\boldsymbol{\mu})=\sum_{k\in\mathcal{K}}\delta_{\mu_{k}}/K$ equal to the empirical distribution of $\boldsymbol{\mu}$ (with $\delta_{\mu_{k}}$ denoting the Dirac measure at $\mu_{k}$):
$\mu^{\prime}\sim M(\boldsymbol{\mu}),\;\;X^{\prime}\mid\mu^{\prime}\sim\mathrm{N}(\mu^{\prime},1).$ (12)
The fundamental theorem of compound decision theory states the following. Given any fixed $s:\mathbb{R}\to\mathbb{R}$, we have that
$\frac{1}{K}\sum_{k\in\mathcal{K}}\mathbb{E}^{\boldsymbol{\mu}}[(s(X_{k})-\mu_{k})^{2}]=\sum_{k\in\mathcal{K}}\frac{1}{K}\mathbb{E}^{\mu_{k}}[(s(X_{k})-\mu_{k})^{2}]=\mathbb{E}^{M(\boldsymbol{\mu})}[(s(X^{\prime})-\mu^{\prime})^{2}].$ (13)
In the left hand side display, $\boldsymbol{\mu}\in\mathbb{R}^{K}$ is treated as fixed (as in (10)), while in the right hand side display, we only have a single random $\mu^{\prime}\in\mathbb{R}$ that is randomly drawn from $M(\boldsymbol{\mu})$ (as in (12)). From (13) it immediately follows that the optimal simple separable estimator is the Bayes estimator under $M(\boldsymbol{\mu})$, that is $s(x)=\mathbb{E}^{M(\boldsymbol{\mu})}[\mu^{\prime}\mid X^{\prime}=x]$. This construction motivated Robbins to construct feasible non-separable estimators $\hat{\mu}_{k}=\hat{\mu}_{k}(X)$ that have risk close to the optimal simple separable estimator. [^4] Thus the optimal simple separable estimator defines an oracle benchmark. Just as in classical decision theory, one often restricts attention to subclasses of estimators, e.g., equivariant or unbiased, Robbins set a benchmark defined by simple separable estimators. What is slightly unconventional here is a fundamental asymmetry: the oracle estimator must be simple separable but has access to the true $\boldsymbol{\mu}$, while the feasible estimator is non-separable but must work without knowledge of $\boldsymbol{\mu}$. Empirical Bayes (Robbins, 1956) ideas work well for this task, for example, Jiang and Zhang (2009) provide sharp guarantees for the performance of a nonparametric empirical Bayes method in mimicking the best simple separable estimator in the Gaussian sequence model in (10) as $K\to\infty$.
7.2 The best simple separable compound e-values
The connection of compound e-values to the fundamental theorem of compound decisions is born out from averaging a suitable statistical criterion, e.g., replacing the criterion $\mathbb{E}^{\mathbb{P}}[E_{k}]\leq 1$ for all $\mathbb{P}\in\mathcal{P}_{k}$ by a corresponding requirement summed over all null $k$. To further clarify the connection, we provide a construction of optimal simple separable compound e-values in sequence models with dominated null marginals via the fundamental theorem of compound decisions.
Suppose that $X=(X_{k}:k\in\mathcal{K})$, where $X_{k}$ takes values in a space $\mathcal{X}$. For any $\mathbb{P}\in\mathcal{P}$, let $\mathbb{P}_{k}$ be the $k$-th marginal of $\mathbb{P}$, that is, the distribution of $X_{k}$ when $(X_{k}:k\in\mathcal{K})\sim\mathbb{P}$. Suppose that for all $k\in\mathcal{K}$, the $k$-th hypothesis $H_{k}$ is specified as a simple point null hypothesis on the $k$-th marginal, i.e., $\mathcal{P}_{k}=\{\mathbb{P}\,:\,\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}\}$ for some prespecified distribution $\mathbb{P}_{k}^{\circ}$ with $\mathrm{d}\nu$-density equal to $p_{k}^{\circ}$, where $\nu$ is a common dominating measure. We assume that there exists $\mathbb{P}\in\mathcal{P}$ such that $\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}$ for all $k\in\mathcal{K}$ (as would be the case, e.g., if the $X_{k}$ are independent and we take =×$\mathbb{P}=\bigtimes_{k\in\mathcal{K}}\mathbb{P}_{k}^{\circ}$).
For $k\in\mathcal{K}$, we let $\mathbb{Q}_{k}^{\circ}$ be distributions on $\mathcal{X}$ with $\mathrm{d}\nu$-densities $q_{k}^{\circ}$. We seek to solve the following optimization problem:
$\underset{s(\cdot)}{\text{maximize}}$ $\frac{1}{K}\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{Q}_{k}^{\circ}}[\log(s(X_{k}))]$ (14)
subject to $s:\mathcal{X}\to[0,\infty]$
[…]
Theorem 7.2 (Optimal simple separable compound e-values).
The optimal solution to optimization problem (14) is given by the likelihood ratio of mixtures over the null and alternative:
$s(x)=\frac{\sum_{j\in\mathcal{K}}q_{j}^{\circ}(x)}{\sum_{j\in\mathcal{K}}p_{j}^{\circ}(x)}.$ (15)
Proof.
Let $\mathbb{P}^{\circ}$ be a probability measure with $k$-th marginal given by $\mathbb{P}_{k}^{\circ}$ for all $k\in\mathcal{K}$. We will first solve the optimization problem subject to the weaker constraint that $s(X_{1}),\dotsc,s(X_{K})$ satisfy $\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{P}^{\circ}}[s(X_{k})]\leq K$.
We will next provide an argument inspired by Robbins’ fundamental theorem of compound decisions. Define the mixture distributions ¯$\overline{\mathbb{P}}^{\circ}=\sum_{k\in\mathcal{K}}\mathbb{P}_{k}^{\circ}/K$ and ¯$\overline{\mathbb{Q}}^{\circ}=\sum_{k\in\mathcal{K}}\mathbb{Q}_{k}^{\circ}/K$. Then, for any fixed $s:\mathcal{X}\to[0,\infty]$, we have the following formal equalities:
$\frac{1}{K}\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{P}_{k}^{\circ}}[\log(s(X_{k}))]=\mathbb{E}^{X^{\prime}\sim\overline{\mathbb{P}}^{\circ}}[\log(s(X^{\prime}))],\;\;\;\frac{1}{K}\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{Q}_{k}^{\circ}}[s(X_{k})]=\mathbb{E}^{X^{\prime}\sim\overline{\mathbb{Q}}^{\circ}}[s(X^{\prime})].$
Thus, it suffices to solve the following optimization problem:
$\underset{s(\cdot)}{\text{maximize}}$ $\mathbb{E}^{X^{\prime}\sim\overline{\mathbb{Q}}^{\circ}}[\log(s(X^{\prime}))]$
subject to $s:\mathcal{X}\to[0,\infty]$
$\mathbb{E}^{X^{\prime}\sim\overline{\mathbb{P}}^{\circ}}[s(X^{\prime})]\leq 1.$
This is a well-known optimization problem (Kelly, 1956) and has solution given by (15). Finally we may verify in that $s$ in (15) indeed yields compound e-values. To this end, write $\overline{q}^{\circ}$ for the $\mathrm{d}\nu$ density of $\overline{\mathbb{Q}}^{\circ}$. Then, we have the following for any $\mathbb{P}$:
$\sum_{k:\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}}\mathbb{E}^{\mathbb{P}}[E_{k}]=\sum_{k:\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}}\int\frac{K\overline{q}^{\circ}(x)}{\sum_{j\in\mathcal{K}}p_{j}^{\circ}(x)}p_{k}^{\circ}(x)\nu(\mathrm{d}x)\leq\int K\overline{q}^{\circ}(x)\nu(\mathrm{d}x)=K,$
and this completes the proof. ∎
Some remarks are in order:
  • The objective $\mathbb{E}[\log(s)]$ that we use in (14) is a long-standing criterion of optimality; see the classic works of Kelly (1956); Breiman (1961); Bell and Cover (1980), and the more recent works of Shafer (2021); Grünwald et al. (2024); Vovk and Wang (2024); Waudby-Smith and Ramdas (2024). This quantity represents the rate of growth of e-processes in a single-hypothesis setting. Other choices of the objective can be used for (14) and they would lead to different forms of the optimal compound e-values.
  • The solution is closely related to the Optimal Discovery Procedure of Storey (2007), which solves a compound Neyman-Pearson optimization problem.
  • The objective value of the optimization problem is given by $\operatorname{KL}(\overline{\mathbb{Q}}^{\circ}\,||\,\overline{\mathbb{P}}^{\circ})$, where $\operatorname{KL}(\cdot\,||\,\cdot)$ denotes the Kullback-Leibler divergence. By convexity, $\operatorname{KL}(\overline{\mathbb{Q}}^{\circ}\,||\,\overline{\mathbb{P}}^{\circ})\leq\sum_{k\in\mathcal{K}}\operatorname{KL}(\mathbb{Q}_{k}^{\circ}\,||\,\mathbb{P}_{k}^{\circ})/K$, which means that the optimal simple separable compound e-value has a worse objective than using the optimal separable e-value for each individual testing problem of $\mathbb{P}_{k}^{\circ}$ vs $\mathbb{Q}_{k}^{\circ}$ given by $E_{k}^{*}=q_{k}^{\circ}(X_{k})/p_{k}^{\circ}(X_{k})$.
  • In view of the above, why is the optimal simple separable compound e-value of interest? The reason is that if $K$ is large (but the individual statistical problems have limited sample sizes), then it may be possible to mimic the simple separable compound e-value without knowledge of the null $\mathbb{P}_{k}^{\circ}$, e.g., via empirical Bayes methods. We believe such a strategy will be fruitful when the hypotheses $\mathcal{P}_{k}$ are composite, see Section 5.3 above and Section 7.3 below for illustrative examples of how this may work. [^5] The construction of Section 5.3 does not directly build on the compound likelihood ratio in (15) but is conceptually related.
We finally note that any tight simple separable compound e-value for testing point nulls $\mathcal{P}_{k}=\{\mathbb{P}\,:\,\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}\}$ must be of the form (15) for some densities $q_{1},\dots,q_{K}$.
Proposition 7.3.
Suppose that $E_{1},\dots,E_{K}$ are tight simple separable compound e-values (with the nulls defined as in the previous parts of this section). Then there exist $\mathrm{d}\nu$-densities $q_{1},\dots,q_{K}$ such that $E_{k}$ may be represented as in (15) on the support of $\sum_{k\in\mathcal{K}}\mathbb{P}_{k}/K$ and it is without loss of generality to take $q_{1}=\dots=q_{K}$.
Proof.
First, by definition of simple and separable, there exists a function $s$ such that $E_{k}=s(X_{k})$. The supremum in the definition of tightness must be attained when $\mathcal{N}(\mathbb{P})=K$, and so:
$\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{P}_{k}}[s(X_{k})]=K.$
Now define $\overline{p}^{\circ}$ as the $\mathrm{d}\nu$ density of $\overline{\mathbb{P}}^{\circ}$, where the latter object is defined as in Theorem 7.2. Then the above may be written as:
$\int s(x)\overline{p}^{\circ}(x)\mathrm{d}\nu(x)=1.$
Thus, if we define $q_{k}(x)=s(x)\overline{p}^{\circ}(x)$ for all $k\in\mathcal{K}$, we find that $q_{k}$ is indeed a $\mathrm{d}\nu$-density and $s(x)$ may be represented as in (15). ∎
7.3 Compound universal inference
We now slightly modify the setting of Section 7.2 to accommodate composite null hypotheses with nuisance parameters. We continue to suppose that $X=(X_{k}:k\in\mathcal{K})$, where $X_{k}$ takes values in a space $\mathcal{X}$ and that the $k$-th hypothesis $H_{k}$ is specified solely in terms of the marginal distribution of $X_{k}$. However, we now suppose that $H_{k}$ is composite and that $k$-th null marginal may be parameterized by a nuisance parameter $\psi_{k}$ that lies in a parameter space $\Psi$. Concretely, $\mathcal{P}_{k}=\{\mathbb{P}:\mathrm{d}\mathbb{P}_{k}/\mathrm{d}\nu=p^{\circ}_{\psi_{k}}\text{ for some }\psi_{k}\in\Psi\}$, where $\{p^{\circ}_{\psi}:\psi\in\Psi\}$ is a family of $\mathrm{d}\nu$-densities on $\mathcal{X}$. We seek to test each composite null $H_{k}$ against the alternative $\{\mathbb{Q}:\mathrm{d}\mathbb{Q}_{k}/\mathrm{d}\nu=q^{\circ}\}$, where $q^{\circ}$ is a further $\mathrm{d}\nu$-density on $\mathcal{X}$ (which we choose to be the same for all $k\in\mathcal{K}$).
In the above setting, universal inference (UI) of Wasserman et al. (2020) provides a flexible and practical way of constructing separable e-values: [^6] Throughout this section we assume that infima are attained.
$E_{k}^{\mathrm{UI}}=\frac{q^{\circ}(X_{k})}{p^{\circ}_{\widehat{\psi}_{k}^{\mathrm{UI}}}(X_{k})},\,\,\,\widehat{\psi}_{k}^{\mathrm{UI}}\in\operatorname*{argmin}\left\{\frac{q^{\circ}(X_{k})}{p^{\circ}_{\psi}(X_{k})}\,:\,\psi\in\Psi\right\}.$ (16)
If we instead seek to construct compound e-values, we propose compound universal inference (CUI) in which we profile over the nuisance parameters appearing in the optimal separable compound e-values of Theorem 7.2:
$E_{k}^{\mathrm{CUI}}=\frac{Kq^{\circ}(X_{k})}{\sum_{j\in\mathcal{K}}p^{\circ}_{\widehat{\psi}_{j}}(X_{k})}\,\text{ for all }\,k\in\mathcal{K},\,\,\,$ (17)
$\text{where}\;\widehat{\psi}_{1},\dots,\widehat{\psi}_{K}\in\operatorname*{argmin}\left\{\sum_{k\in\mathcal{K}}\frac{Kq^{\circ}(X_{k})}{\sum_{j\in\mathcal{K}}p^{\circ}_{\psi_{j}}(X_{k})}\,:\,\psi_{1},\dotsc,\psi_{K}\in\Psi\right\}.$
We also consider a variation of the above that we call empirical Bayes universal inference (EBUI):777 Nguyen and Gupta (2023) use universal inference for an empirical Bayes application. By contrast, here we use ideas from empirical Bayes and universal inference to provide a new construction of compound e-values.
$E_{k}^{\mathrm{EBUI}}=\frac{q^{\circ}(X_{k})}{\int p^{\circ}_{\psi}(X_{k})\mathrm{d}\widehat{G}(\psi)}\,\text{ for all }\,k\in\mathcal{K},\,\,\,$ (18)
$\text{where}\;\widehat{G}\in\operatorname*{argmin}\left\{\sum_{k\in\mathcal{K}}\frac{q^{\circ}(X_{k})}{\int p^{\circ}_{\psi}(X_{k})\mathrm{d}G(\psi)}\,:\,G\text{ distribution supported on }\Psi\right\}.$
We have the following result according to which compound universal inference leads to compound e-values whose sum is at least as large as the universal inference e-values. This highlights potential power gains from considering compound e-values instead of e-values.
Proposition 7.4.
All of the following are compound e-variables for $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$:
  • $E_{1}^{\mathrm{UI}},\dots,E_{K}^{\mathrm{UI}}$,
  • $E_{1}^{\mathrm{CUI}},\dots,E_{K}^{\mathrm{CUI}}$,
  • $E_{1}^{\mathrm{EBUI}},\dots,E_{K}^{\mathrm{EBUI}}$.
Moreover it holds that:
$\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{CUI}}\geq\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{EBUI}}\geq\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{UI}}.$
Proof.
We first prove the inequalities. Let $\widehat{\psi}_{1},\dots,\widehat{\psi}_{K}$ be the minimizers in (17). Then, the distribution $\sum_{k\in\mathcal{K}}\delta_{\widehat{\psi}_{k}}/K$ is feasible for the minimization in (18) and leads to the same objective value, proving the first inequality. Next:
$\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{EBUI}}$ $=\inf_{G}\sum_{k\in\mathcal{K}}\frac{q^{\circ}(X_{k})}{\int p^{\circ}_{\psi}(X_{k})\mathrm{d}G(\psi)}\geq\sum_{k\in\mathcal{K}}\inf_{G}\frac{q^{\circ}(X_{k})}{\int p^{\circ}_{\psi}(X_{k})\mathrm{d}G(\psi)}=\sum_{k\in\mathcal{K}}\inf_{\psi_{k}\in\Psi}\frac{q^{\circ}(X_{k})}{p^{\circ}_{\psi_{k}}(X_{k})}=\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{UI}}.$
Now, take any $\mathbb{P}\in\mathcal{P}$. For any $k$ with $\mathbb{P}\in\mathcal{P}_{k}$, let $\psi_{k}$ be such that $\mathrm{d}\mathbb{P}_{k}/\mathrm{d}\nu=p^{\circ}_{\psi_{k}}$ (which exists by the definition of $\mathcal{P}_{k}$ in this section). For all other $k$, let $\psi_{k}\in\Psi$ be arbitrary. Then by construction,
$\sum_{k\in\mathcal{K}}\frac{Kq^{\circ}(X_{k})}{\sum_{j\in\mathcal{K}}p^{\circ}_{\psi_{j}}(X_{k})}\geq\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{CUI}}.$
Arguing as in the proof of Theorem 7.2, the summands in the left hand side of the above inequality are compound e-values. Hence $E_{1}^{\mathrm{CUI}},\dots,E_{K}^{\mathrm{CUI}}$ are compound e-values (and similarly for EBUI and UI). ∎
8 Summary
We offer a formal treatment of asymptotic and compound e-values by giving them rigorous definitions. We provide several prototypical constructions, and explain why compound e-values are a fundamental concept in multiple testing. We demonstrate several ways of constructing approximate or exact compound e-values, for example,
  1. 1.
    by leveraging any FDR controlling procedure,
  2. 2.
    through the compound likelihood ratio (a ratio of mixture likelihoods) which yields log-optimal simple separable compound e-values,
  3. 3.
    through compound universal inference,
  4. 4.
    or by calibration of average significance controlling p-values.
Since notions related to compound e-values have already been extensively used in the recent multiple testing literature, we believe that our formalism will help streamline the development of new procedures, and provide a common language for future research in multiple testing that builds upon compound e-values.
We briefly point out that the term compound p-values has previously been used in a way that is distinct from Definition 4.1. Habiger and Peña (2014) considered a setting akin to the sequence model in Section 7 wherein $X=(X_{k}:k\in\mathcal{K})$ and in principle $X_{k}$ can be used to test $H_{k}$ for all $k\in\mathcal{K}$. Let $P_{1},\dots,P_{K}$ be $[0,1]$-valued random variables. Habiger and Peña (2014) called these “compound p-values” if they are bona fide p-values (that is $\mathbb{P}(P_{k}\leq t)\leq t$ for all $t\in[0,1]$ and all $\mathbb{P}\in\mathcal{P}_{k}$) and depend also on data for other hypotheses $(X_{\ell}:\ell\neq k)$, i.e., $P_{k}=P_{k}(X)$ rather than $P_{k}=P_{k}(X_{k})$. In other words, “compound p-values” are not separable as per Definition 7.1, and so could be called “non-separable p-values” instead. The definition treats separable p-values as the default data analysis choice (which is reasonable given how p-values are often used in practice), and one needs a qualifier to refer to non-separable p-values. By contrast, we prefer to think of (compound) p-values and e-values as being non-separable by default (given their flexibility and generality), and to use an explicit qualifier for the separable case. Moreover, we use the term compound in relation to the averaging (compounding) operation inherent in Definition 4.1.
Appendix A Compound e-values as weights in merging p-values
A vector of arbitrarily dependent p-variables can be merged into one p-variables through p-merging functions in the sense of Vovk and Wang [2020]. All admissible homogeneous p-merging functions can be written as
$F(\mathbf{p})=\inf\left\{\alpha\in(0,1):\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{p_{k}}{\alpha}\right)\geq 1\right\}\mbox{ ~{}~{}~{}for $\mathbf{p}=(p_{1},\dots,p_{K})\in[0,\infty)^{K}$,}$ (19)
for some $\lambda_{1},\dots,\lambda_{K}$ that are nonnegative weights summing to $1$ and admissible p-to-e calibrators $f_{1},\dots,f_{K}$ (Vovk et al. [2022, Theorem 5.1]). Here, we set $\inf\emptyset=1$.
Consider the global null setting $\mathbb{P}\in\bigcap_{k\in\mathcal{K}}\mathcal{P}_{k}$, and let $E_{1},\dots,E_{K}$ be compound e-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$. Write $\mathbf{E}=(E_{1},\dots,E_{K})$.
We follow the setting of Ignatiadis et al. [2024], where one has access to a vector of p-values for the hypotheses $H_{1},\dots,H_{K}$ that is independent of $\mathbf{E}$. This may be relevant in the context of follow-up experiments. Suppose that $\mathbf{P}$ is a vector of p-variables for $\mathbb{P}$ independent of $\mathbf{E}$. We can define the e-weighted version of (19) as
$F(\mathbf{p},\mathbf{E})=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}E_{k}f_{k}\left(\frac{p_{k}}{\alpha}\right)\geq 1\right\}\mbox{ ~{}~{}~{}for $\mathbf{p}=(p_{1},\dots,p_{K})\in[0,\infty)^{K}$.}$ (20)
The formulation (20) includes (19) as a special case via $E_{k}=\lambda_{k}K$ for $k\in\mathcal{K}$. Note that for any calibrators $f_{1},\dots,f_{K}$, the function $F(\mathbf{p},\mathbf{e})$ is decreasing in $\mathbf{e}$. This is intuitive, as larger e-values carry evidence against the null hypothesis, in the same direction as smaller p-values.
For a clear comparison with some classic p-merging functions, in what follows, we allow $F$ in (20) to be defined with any decreasing functions $f_{1},\dots,f_{K}$ satisfying $\int_{0}^{1}f_{k}(x)\mathrm{d}x\leq 1$, $f_{k}\geq 0$ on $[0,1]$, and $f_{k}\leq 0$ on $(1,\infty)$ for each $k\in\mathcal{K}$. These functions are more general than p-to-e calibrators, and they generate p-merging functions dominated by those generated by calibrators.
Proposition A.1.
In the above setting, we have
$\mathbb{P}(F(\mathbf{P},\mathbf{E})\leq\alpha)\leq\alpha~{}~{}~{}\mbox{for all$\alpha\in(0,1)$}.$
Proof.
Let $\beta\in(0,1)$. Note that $f_{k}^{+}:=f_{k}\vee 0$ is a calibrator. For any calibrator $f$ and p-variable $P$, we can check that $\beta^{-1}f({P}/{\beta})$ is an e-variable for $\mathbb{P}$; see e.g., Lemma 2.3 of Gasparin et al. [2025]. Therefore,
$\mathbb{E}^{\mathbb{P}}\left[\frac{1}{K}\sum_{k=1}^{K}E_{k}f_{k}^{+}\left(\frac{P_{k}}{\beta}\right)\right]\leq\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}^{\mathbb{P}}[E_{k}]\beta\leq\beta.$
Markov’s inequality gives
$\mathbb{P}\left(\frac{1}{K}\sum_{k=1}^{K}E_{k}f_{k}\left(\frac{P_{k}}{\beta}\right)\geq 1\right)\leq\mathbb{P}\left(\frac{1}{K}\sum_{k=1}^{K}E_{k}f_{k}^{+}\left(\frac{P_{k}}{\beta}\right)\geq 1\right)\leq\beta.$
Using the monotonicity of $\beta\mapsto f_{k}(p_{k}/\beta)$, we have, for any $\alpha\in(0,1)$,
$F(\mathbf{P},\mathbf{E})\leq\alpha\iff\frac{1}{K}\sum_{k=1}^{K}E_{k}f_{k}\left(\frac{p_{k}}{\beta}\right)\geq 1\mbox{~{}for all $\beta>\alpha$}.$
Therefore,
$\mathbb{P}(F(\mathbf{P},\mathbf{E})\leq\alpha)$ $=\mathbb{P}\left(\bigcap_{\beta>\alpha}\left\{\frac{1}{K}\sum_{k=1}^{K}E_{k}f_{k}\left(\frac{p_{k}}{\beta}\right)\geq 1\right\}\right)$
$\leq\inf_{\beta>\alpha}\mathbb{P}\left(\frac{1}{K}\sum_{k=1}^{K}E_{k}f_{k}\left(\frac{p_{k}}{\beta}\right)\geq 1\right)\leq\inf_{\beta>\alpha}\beta=\alpha,$
showing the desired result. ∎
Some examples are discussed below. We will see that negative values for $f$ on $(1,\infty)$ allow for recovering some classic p-merging functions.
  1. (i)
    Take $f_{1},\dots,f_{K}$ as $f(x)=2-2x$. We have
    $F(\mathbf{p},\mathbf{E})=\inf\left\{\alpha\in(0,1):\frac{2}{K}\sum_{k=1}^{K}E_{k}(1-P_{k}/\alpha)\geq 1\right\}=\left(\frac{\sum_{k=1}^{K}E_{k}P_{k}}{(\sum_{k=1}^{K}E_{k}-K/2)_{+}}\right)\wedge 1,$
    where we can set $F(\mathbf{p},\mathbf{E})=1$ if the denominator is 00. This can be seen as a e-weighted version of “twice the mean” (truncated at $1$), which is precisely the case if $E_{1}=\dots=E_{K}=1$.
  2. (ii)
    Take $f_{1},\dots,f_{K}$ as $f(x)=-\log x$. By writing $E=\sum_{k=1}^{K}E_{k}/K$, we have
    $F(\mathbf{p},\mathbf{E})=\inf\left\{\alpha\in(0,1):\frac{-1}{K}\sum_{k=1}^{K}E_{k}\log(P_{k}/\alpha)\geq 1\right\}=\left(\exp(E)\left(\prod_{k=1}^{K}P_{k}^{E_{k}}\right)^{1/(KE)}\right)\wedge 1.$
    This can be seen as a e-weighted version of “$\exp(1)$ times the geometric mean”(truncated at $1$), which is precisely the case if $E_{1}=\dots=E_{K}=1$.
There is also a similar e-weighted version of the harmonic mean, but its formula is more involved, and we omit it here.